### v1

# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

Abstract.
We study the sine–Gordon measure defined on each homotopy class. The energy space decomposes into infinitely many such classes indexed by the topological degree $Q\in\mathbb{Z}$. Even though the sine–Gordon action admits no minimizer in homotopy classes with $|Q|\geq 2$, we prove that the Gibbs measure on each class nevertheless concentrates and exhibits Ornstein–Uhlenbeck fluctuations near the multi-soliton manifold in the joint low-temperature and infinite-volume limit. Furthermore, we show that the joint distribution of the multi-soliton centers coincides with the ordered statistics of independent uniform random variables, so that each soliton’s location follows a Beta distribution.
Key words and phrases:
topological solitons; multi-soliton manifold; Ornstein–Uhlenbeck fluctuations, Large deviations, expected location, expected gap
2020 Mathematics Subject Classification:
60F10, 60F05, 82B21, 60H30
1. Introduction
1.1. Motivations and implications of the main results
The sine–Gordon model plays a central role as a fundamental example of a nonlinear scalar field theory admitting topological solitons. In this paper, we specifically study the massless sine–Gordon field theory, with action
$E(\phi)=\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\phi|^{2}dx+\int_{\mathbb{R}}(1-\cos\phi)dx.$ (1.1)
This field theory allows the energy space to be classified into infinitely many disjoint homotopy classes according to the topological invariant $Q$, defined in (1.2). For any finite-energy configuration $E(\phi)<\infty$, the field must satisfy
$\phi(\infty):=\lim_{x\to\infty}\phi(x)\in 2\pi\mathbb{Z},\qquad\phi(-\infty):=\lim_{x\to-\infty}\phi(x)\in 2\pi\mathbb{Z}.$
These boundary conditions imply that the map $x\mapsto e^{i\phi(x)}$ winds around the target circle an integer number of times as $x$ runs from -$\infty$ to $\infty$. This integer $Q$ defines the topological degree/charge or winding number
$Q=\frac{1}{2\pi}(\phi(\infty)-\phi(-\infty))\in\mathbb{Z}.$ (1.2)
According to the winding number $Q$, the energy space $\mathcal{C}$ decomposes into disjoint connected components, referred to as homotopy classes or topological sectors $\mathcal{C}_{Q}$
$\mathcal{C}:=\{\phi\in H_{\textup{loc}}^{1}(\mathbb{R}):E(\phi)<\infty\}=\bigsqcup_{Q\in\mathbb{Z}}\mathcal{C}_{Q},$
where
$\mathcal{C}_{Q}=\{\phi\in\mathcal{C}:(\phi(\infty)-\phi(-\infty))/2\pi=Q\}.$ (1.3)
Within each sector $\mathcal{C}_{Q}$, fields can be continuously deformed into one another. However, a configuration in $\mathcal{C}_{Q}$ cannot be continuously deformed into a configuration in $\mathcal{C}_{Q^{\prime}}$ when $Q\neq Q^{\prime}$. Therefore, when studying the minimization problem for the energy functional $E(\phi)$, we fix a topological sector $\mathcal{C_{Q}}$ and consider
$\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
The following facts are well known:
  • (i)
    $Q=0$ (vacuum sector): the minimizers are the vacuum states
    $\phi(x)=2\pi k,\quad k\in\mathbb{Z}.$
  • (ii)
    $|Q|=1$ (kink/antikink sector): the minimizers are the kinks when $Q=1$ and the anti-kinks when Q=1Q=-1, unique up to translation symmetry
    $\{m(\cdot-\xi)\}_{\xi\in\mathbb{R}}\quad\textup{and}\quad\{m^{-}(\cdot-\xi)\}_{\xi\in\mathbb{R}},$
    where
    $m(x)=4\arctan(e^{x})\quad\textup{and}\quad m^{-}(x)=4\arctan(e^{-x}).$
    Thus, the family of minimizers forms a one-dimensional soliton manifold. These topological solitons, kinks and antikinks, interpolate between the vacua 0 and $2\pi$.
  • (iii)
    $|Q|\geq 2$ (higher-charge sectors): no minimizer exists. For higher charge, the lack of compactness prevents existence of a minimizer.
Although the higher-charge sectors $|Q|\geq 2$ admit no minimizer, we nevertheless study, for each $Q\in\mathbb{Z}$, the concentration and fluctuation behavior of the Gibbs measure [^1] Here $Z_{\varepsilon}$ denotes partition functions that may differ from line to line. $\rho_{\varepsilon}^{Q}$ on the homotopy class $\mathcal{C}_{Q}$
$\rho_{\varepsilon}^{Q}(d\phi)=Z_{\varepsilon}^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx-\frac{1}{2\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x)$
in the joint low-temperature $\varepsilon\to 0$ and infinite-volume $L_{\varepsilon}\to\infty$ limits. For the precise definition of the Gibbs measure $\rho_{\varepsilon}^{Q}$ on each homotopy class $\mathcal{C}_{Q}$, see the next subsection.
We first state our main results in a somewhat informal manner; see Theorems 1.2, 1.4, and 1.6 for the precise statements. In the following, we only consider the nontrivial topological sector $Q\neq 0$, where solitons appear.
Theorem 1.1.
Let $Q\in\mathbb{Z}$ with $Q\neq 0$.
  • (1)
    Under the ensemble $\rho^{Q}_{\varepsilon}$ with $Q>0$, the field $\phi$ exhibits the typical behavior
    $\phi(x)\approx\sum_{j=1}^{Q}m(x-\xi_{j})+\varepsilon^{\frac{1}{2}}\cdot\textup{Ornstein–Uhlenbeck}$
    with
    $\min_{i\neq j}|\xi_{j}-\xi_{i}|\geq\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|\to\infty$
    as $\varepsilon\to 0$ and $L_{\varepsilon}\to\infty$. When $Q<0$ the soliton $m$ is replaced by $m^{-}$.
  • (2)
    As $\varepsilon\to 0$, the joint distribution of the centers $(\xi_{1},\dots,\xi_{|Q|})$ is the ordered statistics of $|Q|$ independent uniform random variables. In particular, each marginal $\xi_{(j)}$ has a Beta distribution. Consequently, the expected position and the expected gap are
    $\mathbb{E}[\xi_{(j)}]\approx-L_{\varepsilon}+\frac{2L_{\varepsilon}j}{|Q|+1},\qquad\mathbb{E}[\xi_{(j)}-\xi_{(j-1)}]\approx\frac{2L_{\varepsilon}}{|Q|+1},$
    where $\xi_{(j)}$ denotes the $j$-th ordered center in increasing rearrangement $\xi_{(1)}\leq\cdots\leq\xi_{(|Q|)}$.
The main results have the following implications:
  • (i)
    The main theorem provides the first result on the concentration and fluctuation behavior of Gibbs measures around multi-solitons $\sum_{j=1}^{Q}m(\cdot-\xi_{j})$, rather than a single soliton. In particular, our result shows that even though the higher-charge sector $\mathcal{C}_{Q}$, $|Q|\geq 2$, does not contain minimizers of the action on $\mathbb{R}$, the Gibbs measure over $\mathcal{C}_{Q}$ still exhibits concentration and fluctuation around multi-soliton configurations.
  • (ii)
    At leading order, typical configurations under the Gibbs ensemble consist of exactly $|Q|$ solitons whose mutual separations are of order $\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$. Consequently, configurations in which the solitons collide are unlikely, and the solitons behave as effectively non-interacting objects.
  • (iii)
    Our base measure is the Brownian bridge (1.4) without a mass term, which lacks correlation decay, whereas the fluctuation measure is the Ornstein–Uhlenbeck measure, exhibiting strong correlation decay. This contrast is rather striking, since in many quantum field and statistical physics models, the fluctuations are typically governed by the underlying base field, rather than having a completely different covariance structure.
  • (iv)
    The expected soliton centers $\xi_{1},\dots,\xi_{|Q|}$ are evenly spaced, dividing the interval $L_{\varepsilon},L_{\varepsilon}]$ into $|Q|+1$ equal parts of length $\tfrac{2L_{\varepsilon}}{|Q|+1}$. Furthermore, each individual center $\xi_{j}$ exhibits a Beta-type fluctuation around its expected position.
Figure 1. a multi-soliton $\sum_{j=1}^{Q}m(\cdot-\xi_{j})$ with $Q=3$ and $(\xi_{1},\xi_{2},\xi_{3})=(-5,10,20)$
The remarkable point is that in previous works (see Subsections 1.3.2, 1.3.3, 1.3.4, and 1.3.5), the behavior of Gibbs measures was mainly studied around a single soliton, and most of the analysis in those works focused primarily on the concentration of the measure. On the other hand, our result is the first study of the Gibbs measure around multi-soliton configurations, involving not only concentration but also a central limit theorem behavior around the multi-soliton manifold. Furthermore, we provide a concrete description of the soliton locations and gaps.
Moreover, the geometry of the multi-soliton manifold $\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):\xi_{j}\in\mathbb{R}\}$ is a central object of interest in this work. In contrast to the single-soliton manifold $\{m(\cdot-\xi):\xi\in\mathbb{R}\}$, the multi-soliton manifold fails to be differentiable and becomes merely a topological manifold when solitons collide, that is, when $|\xi_{i}-\xi_{j}|\lesssim 1$. As a result, fundamental geometric objects such as tangent and normal vectors are no longer well defined in the collision regime. This geometric degeneracy requires a careful analysis and a delicate decomposition of the multi-soliton manifold. See Remark 4.1.
We finally remark that, unlike in the one-dimensional setting where topological solitons are well understood (see Subsection 1.3.1), the situation in two dimensions is quite different. The sine–Gordon equation in 2D also admits soliton-like solutions, often called kink walls, obtained by extending the one-dimensional kink uniformly in another spatial direction. These configurations solve the equation but necessarily have infinite energy, and thus are not finite-energy solitons in the usual sense. To the best of our knowledge, such infinite-energy kink-wall solutions are far less understood and have not been studied as systematically as their one-dimensional counterparts.
1.2. Main results
In this subsection, we present the three main theorems, 1.2, 1.4, and 1.6. Before stating the theorems, we first study Gibbs measures corresponding to each topological degree $Q\in\mathbb{Z}$.
Based on the definition of the topological degree $Q$ in (1.2), when $\phi(\infty)=2\pi n^{+}$ and $\phi(-\infty)=2\pi n^{-}$, with $n^{+},n^{-}\in\mathbb{Z}$, the homotopy class $\mathcal{C}_{Q}$ depends only on the difference
$Q=n_{+}-n_{-},$
not on the individual values of $n_{+}$ and $n_{-}$. Therefore, when describing each sector $\mathcal{C}_{Q}$ with $Q>0$, we fix a representative in the equivalence class by choosing the left boundary value $\phi(-L_{\varepsilon})=0$ as the base point, so that $\phi(L_{\varepsilon})=2\pi Q$, where $L_{\varepsilon}\to\infty$. A similar convention applies for $Q<0$ by reversing the orientation. By symmetry, we only consider the case $Q>0$ throughout the paper, unless specified otherwise.
We now introduce the base measure, namely the Brownian bridge,
$\mu_{\varepsilon}^{Q}(d\phi)=\frac{1}{Z_{\varepsilon}^{Q}}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}d\phi(x),$ (1.4)
which is the Gaussian measure conditioned on $\phi(-L_{\varepsilon})=0$ and $\phi(L_{\varepsilon})=2\pi Q$. This Gaussian measure describes fluctuations around the affine line connecting the boundary values 0 and $2\pi Q$. See Subsection 2.2. This choice of base measure pins down a representative within the equivalence class $\mathcal{C}_{Q}$ and yields a unique Brownian bridge measure $\mu^{Q}_{\varepsilon}$.
For each topological degree $Q\in\mathbb{Z}$, we now define the Gibbs measure, using the Brownian bridge
$\rho_{\varepsilon}^{Q}(d\phi)=Z_{\varepsilon}^{-1}\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi(x))dx\Big\}\mu_{\varepsilon}^{Q}(d\phi).$ (1.5)
In the following, any field $\phi$ distributed according to the Gibbs measure $\rho_{\varepsilon}^{Q}$ is viewed as a function on $\mathbb{R}$, extended trivially by 0 and $2\pi|Q|$ outside the interval $L_{\varepsilon},L_{\varepsilon}]$. We now state, for each $Q\in\mathbb{Z}$, how the Gibbs measure $\rho^{Q}_{\varepsilon}$, associated with the homotopy class $\mathcal{C}_{Q}$, concentrates around the multi-soliton manifold.
Theorem 1.2.
Let $Q\in\mathbb{Z}$ with $Q\neq 0$, and $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ with $\eta>0$ arbitrarily small but fixed.
  • (i)
    There exists $c>0$ such that for any $\delta>0$
    $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2},$ (1.6)
    where dist denotes the $L^{2}(\mathbb{R})$-distance and the multi-soliton manifold [^2] The multi-soliton profile $m(\cdot-\xi_{1})+\dots+m(\cdot-\xi_{Q})$ is invariant under permutation of the labels $i=1,\ldots,|Q|$. Each unordered configuration corresponds to $Q!$ identical ordered configurations. Because of the indistinguishable nature, we work with the ordered set. is defined as
    $\mathcal{M}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\Big\},\quad Q>0.$ (1.7)
    For $Q<0$, we define the corresponding multi-soliton manifold with anti-kinks.
  • (ii)
    There exists $c>0$ such that
    $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d},$ (1.8)
    for any $\delta>0$, $d>0$ satisfying [^3] Later, we prove that the collision manifold $\mathcal{M}_{Q}^{<d}$ is an unlikely event. Therefore, when the field $\phi$ is sufficiently close to $\mathcal{M}_{Q}^{<d}$, in the sense that $\frac{1}{2c}e^{-d}\geq\delta$, we obtain the same result. See Lemmas 3.10 and 3.12. $\frac{1}{2c}e^{-d}\geq\delta$, where the collison manifold $\mathcal{M}_{Q}^{<d}$, $Q>0$, is defined as
    $\mathcal{M}^{<d}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\}.$ (1.9)
    For $Q<0$, we define the corresponding manifold with anti-kinks.
The first part of Theorem 1.2 shows that when a field $\phi$ is far from the multi-soliton manifold $\mathcal{M}_{Q}$, the event is exponentially unlikely with rate $\delta^{2}$. In particular, the manifold consists of exactly $|Q|$ solitons, without any mixture of kinks and anti-kinks. Furthermore, the second part shows that although a field $\phi$ is close to the multi-soliton manifold, when the solitons collide so that $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, the collision region becomes a large-deviation event. From (1.6), (1.8), and the condition $\frac{1}{2c}e^{-d}\geq\delta$, we may choose the distance and collision scales as $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$. With these choices, we obtain
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta_{\varepsilon}\}\big)\leq e^{-c\log\frac{1}{\varepsilon}}$
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\leq e^{-c\varepsilon^{-\frac{1}{2}+}}$
for some $c>0$ as $\varepsilon\to 0$. An interesting fact is that we identify the collision scale $\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ that determines the typical behavior of solitons. Hence, most of the probability mass is concentrated in the well-separated (non-collision) region
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\geq d_{\varepsilon}})<\delta_{\varepsilon}\},$
where the non-collision manifold $\mathcal{M}^{\geq d_{\varepsilon}}_{Q}$ is defined as
$\mathcal{M}^{\geq d_{\varepsilon}}_{Q}:=\Big\{\sum_{j=1}^{Q}m(\cdot-\xi_{j}):-\infty<\xi_{1}\leq\dots\leq\xi_{Q}<\infty\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\Big\}.$
In the proof of Theorem 1.2, the key aspect is to understand how the solitons interact with each other on the collision scale $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, and how the energy behaves even though these configurations are not minimizers.
Remark 1.3.
In Theorem 1.2, there is a competition between the vanishing energy scale $\varepsilon\to 0$ and the entropic effects arising from the growing interval $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}\to\infty$. Under our method, the interval size $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ is optimal. See Remark 6.2 for an explanation of this scaling.
As a consequence of Theorem 1.2, under the measure $\rho^{Q}_{\varepsilon}$, the leading-order behavior is described by $\sum_{j=1}^{Q}m(\cdot-\xi_{j})$ with almost no collision $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}|\to\infty$. In the following theorem, we investigate the next-order fluctuation behavior around the multi-solitons. To state our next theorem, we first introduce $\pi^{\varepsilon}$, the projection onto the (approximating) [^4] Since we work on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. As $\varepsilon\to 0$, $m^{\varepsilon}(\cdot-\xi_{j})$ becomes a more and more precise approximation of the topological soliton $m(\cdot-\xi_{j})$ on $\mathbb{R}$ (see (4.3)) multi-soliton manifold $\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d_{\varepsilon}}$ defined in (4.15), where the measure $\rho^{Q}_{\varepsilon}$ concentrates (see Remarks 6.4, 6.5, and (6.32)).
Theorem 1.4.
Let $Q\in\mathbb{Z}$ with $Q\neq 0$, $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$, and let $F$ be a bounded and continuous function. Then
$\lim_{\varepsilon\to 0}\int F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\rho_{\varepsilon}^{Q}(d\phi)=\int F(\phi)\mu_{\text{OU}}(d\phi),$
where $\mu_{\text{OU}}$ is the Ornstein–Uhlenbeck measure
$\mu_{\text{OU}}(d\phi)=Z^{-1}\exp\Big\{-\frac{1}{2}\langle\phi,(-\partial_{x}^{2}+1)\phi\rangle_{L^{2}(\mathbb{R})}\Big\}\prod_{x\in\mathbb{R}}d\phi(x).$
The fluctuations described in Theorem 1.4 exhibit behavior that is different from the classical result of Ellis–Rosen [19, Theorem 4], where Ellis–Rosen studied the central limit theorem for general Gibbs measures in the low–temperature limit. In the usual low-temperature setting $\varepsilon\to 0$, the fluctuation behavior is determined by the second variation $\nabla^{2}E$ of the energy evaluated at the minimizers. In contrast, Theorem 1.4 concerns a joint limit in which $\varepsilon\to 0$ and $L_{\varepsilon}\to\infty$ simultaneously. This introduces a competition between energy and entropic effects, leading to a fluctuation behavior that differs markedly from that of Ellis–Rosen [19, Theorem 4]. Furthermore, in our case, the energy $E$ does not possess minimizers in the homotopy classes with $|Q|\geq 2$. Consequently, the second–variation approach used in Ellis–Rosen [19, Theorem 4] cannot be applied here.
To prove Theorem 1.4, although a minimizer does not exist when $|Q|\geq 2$, we analyze the second variation of the energy at the multi–soliton configuration $m_{\xi_{1},\dots,\xi_{Q}}=\sum_{j=1}^{Q}m(\cdot-\xi_{j})$
$\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})=-\partial_{x}^{2}+\cos(m_{\xi_{1},\dots,\xi_{Q}})$
under the separation scale $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ (see also (5.1)). It allows to study the Gaussian measure $\exp\big\{-\frac{1}{2}\langle\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})v,v\rangle\big\}$ whose covariance structure (Lemma 9.3) and correlation decay (Proposition 8.3) are crucial ingredients in the proof of the central limit theorem.
In particular, the spectral analysis of $\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})$ is closely linked to an understanding of the geometry of the multi-soliton manifold $\mathcal{M}_{Q}$ through quadratic forms such as
$\langle\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})v,v\rangle,$
where $v$ is taken in either the tangential or the normal direction to the manifold $\mathcal{M}_{Q}$. As discussed in Remark 4.1, the multi-soliton manifold $\mathcal{M}_{Q}$ and, in particular, the collision manifold $\mathcal{M}_{Q}^{<d}$ fail to be differentiable and are only a topological manifold. This lack of smooth geometric structure prevents the use of standard tools such as tangent and normal decompositions, which are essential for performing a second-order expansion and identifying Gaussian fluctuations in Ellis–Rosen [19, Theorem 4]. To overcome this issue, by proving the large-deviation theorem 1.4, we exclude the collision manifold $\mathcal{M}_{Q}^{<d}$. On the resulting non-collision manifold $\mathcal{M}_{Q}^{\geq d}$, the manifold is smooth and admits well-defined tangent and normal directions. This allows us to carry out a geometric decomposition for studying $\nabla^{2}E(m_{\xi_{1},\dots,\xi_{Q}})$.
Remark 1.5.
In Theorem 1.4, the restriction $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ follows from Theorem 1.2. If one could enlarge the admissible range of $L_{\varepsilon}$ in Theorem 1.2, then the fluctuation result in Theorem 1.4 continues to hold on a much larger scale.
We now state the final theorem. The infinite separation condition $|\xi_{i}-\xi_{j}|\to\infty$ implies that the interactions between solitons are negligible. However, it does not provide any information about the locations of the solitons. In the following, we analyze the joint and marginal distribution of the soliton locations $(\xi_{1},\dots,\xi_{|Q|})$, which describes their expected positions and the gaps.
Before stating the final theorem, we first present some preliminaries. Under the coordinate representation [^5] Thanks to the large deviation results in Theorem 1.2, we can write the field $\phi$ as a multi–soliton configuration plus a small perturbation. $\phi=\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})+v$, where $\|v\|_{L^{2}}<\delta$, defined in (4.15), we consider the projection $\pi_{\varepsilon}^{T}(\phi)=(\xi_{1},\dots,\xi_{Q})$ onto the coordinate variables, where666when working on the finite volume $L_{\varepsilon},L_{\varepsilon}]$, we need to define an approximating multi–soliton profile $\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j})$ so that the entire transition from 0 to $2\pi Q$ occurs inside this interval. For this reason, we introduce a slightly smaller interval, denoted by $\overline{L}_{\varepsilon}\sim L_{\varepsilon}$, on which the translation $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ modes are defined (see (4.2)). $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$. Furthermore, we define the (marginal) tangential projection $\pi^{T}_{j}(\phi)=\xi_{j}$, where $\xi_{j}$ denotes the $j$-th ordered center in the increasing rearrangement $\xi_{1}\leq\cdots\leq\xi_{Q}$.
Theorem 1.6.
Let $Q\in\mathbb{Z}$ with $Q\neq 0$, $\overline{L}_{\varepsilon}\sim L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$.
  • (i)
    The joint distribution of the centers $(\xi_{1},\dots,\xi_{|Q|})$ is given by the ordered statistics of $|Q|$ independent uniform random variables on $\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$
    $\rho_{\varepsilon}^{Q}\{\pi_{\varepsilon}^{T}(\phi)\in A\}=\frac{\big|A\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{|Q|}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{|Q|}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+}))$
    as $\varepsilon\to 0$, where $A\subset\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$ is a measurable subset.
  • (ii)
    The marginal distribution, that is, each center $\xi_{j}$, has a Beta-shaped fluctuation
    $\rho_{\varepsilon}^{Q}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\int_{B}f_{j}(x)dx\cdot(1+O(\varepsilon^{0+}))$
    as $\varepsilon\to 0$, where $B\subset[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ is a measurable subset and
    $f_{j}(x)=\frac{|Q|!}{(2L_{\varepsilon})^{|Q|}}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{|Q|-j}}{(|Q|-j)!},\;\;-\overline{L}_{\varepsilon}<x<\overline{L}_{\varepsilon}.$
  • (iii)
    The expected location of each soliton is given by
    $\mathbb{E}_{\rho_{\varepsilon}^{Q}}\big[\pi^{T}_{j}(\phi)\big]=\Big(-\overline{L}_{\varepsilon}+\frac{2\overline{L}_{\varepsilon}j}{|Q|+1}\Big)\cdot(1+O(\varepsilon^{0+})).$
In Theorem 1.6, we obtain explicit formulas for the joint and marginal distributions, thereby describing quantitatively how the solitons are arranged over the entire interval. In particular, Theorem 1.6 shows that the expected centers $\xi_{j}$ are evenly spaced, dividing the interval $\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ into $|Q|+1$ equal parts, each of length $\frac{2\overline{L}_{\varepsilon}}{|Q|+1}$.
Notice that, in contrast with the soliton resolution result of Chen–Liu–Lu [14, Theorem 1.1], which describes the asymptotic decomposition of solutions under deterministic dynamics, our results are probabilistic in nature. Under the Gibbs measure, we identify the typical locations of individual solitons and precisely characterize the gaps between neighboring solitons.
1.3. Related literature
1.3.1. Sine–Gordon field theory
McKean-Vaninsky [32] studied the construction of the one–dimensional sine Gordon measure. More recently, Lacoin–Rhodes–Vargas [27] studied the one–dimensional sine–Gordon measure with a log–correlated base field in the full subcritical regime on a bounded domain.
From the PDE perspective, the one-dimensional sine–Gordon equation and its soliton solutions have been widely studied in recent years. In particular, the asymptotic stability and long–time behavior of soliton solutions have attracted considerable attention. We refer to the works of McKean [30], Lührmann–Schlag [28], Alejo–Muñoz–Palacios [1], Chen–Liu–Lu [14], and Chen–Lührmann [15].
The two-dimensional sine–Gordon theory has connections to various problems in statistical physics, such as the Coulomb gas and the XY model. We refer to [21, 10, 18, 4, 5, 3, 23] for the study of the 2D sine–Gordon measures and to [25, 11, 9] for the two-dimensional sine–Gordon equation with stochastic forcing. In two dimensions, the sine–Gordon equation also admits soliton-like solutions, often called kink walls, obtained by extending the one-dimensional kink uniformly in another spatial direction. Although these configurations solve the 2D sine–Gordon equation, they necessarily have infinite energy and hence fall outside the standard class of finite-energy solitons (see [29, Chapter 5.4]). To the best of our knowledge, these infinite-energy kink-wall solutions are considerably less explored and have not been developed to the same extent as their one-dimensional counterparts.
1.3.2. Continuum focusing NLS Gibbs measure and single-soliton manifold
The behavior of (invariant) Gibbs measures around a single soliton manifold has been studied for the focusing nonlinear Schrödinger (NLS) equation. For the continuum focusing NLS Gibbs measure, McKean [31] initiated the study of the infinite-volume limit. Later, Rider [35] and Tolomeo–Weber [38] proved that, on the 1D torus, the measure concentrates around the single-soliton manifold in the infinite-volume limit. In particular, Tolomeo–Weber [38] identified a critical regime for the strength of the coupling: either the measure strongly concentrates around the single-soliton manifold, or the limiting measure reduces to the underlying Gaussian field. Recently, in [36], the first and third authors proved a central limit theorem for the Gibbs measure around a single-soliton manifold. In this paper we establish the corresponding result for multi-soliton manifolds in the sine-Gordon setting. The two situations exhibit different fluctuation behavior: For the focusing NLS Gibbs measure, the fluctuations are white noise near a single soliton, but in the sine–Gordon case studied here, the system exhibits Ornstein–Uhlenbeck fluctuations near the multi-soliton manifold.
1.3.3. Discrete focusing NLS Gibbs measure and single-soliton manifold
The discretized focusing Gibbs measure and the nonlinear Schrödinger (NLS) equation on a lattice have been studied. When $d\geq 3$, Chatterjee–Kirkpatrick [12] initiated the study of the discretized focusing NLS Gibbs measure, identified a critical temperature, and showed that below this threshold the system exhibits striking single-soliton–like behavior. In [13] Chatterjee used microcanonical invariant measures and showed that a typical function in the ensemble decomposes into a “visible” part, which is close to a single soliton, and an “invisible” part that is small in the $L^{\infty}$ norm. Notice that, in particular, regarding the reason why a single-soliton profile appears, Chatterjee mentioned in [13] that “ whereas multisoliton solutions eventually merge into a single soliton on the finite discrete torus considered in Theorem 1”. In contrast to that situation, in our sine–Gordon model, the boundary conditions imposed by a fixed homotopy class (1.3) enforce the presence of multiple solitons, and therefore prevent the multi-soliton configuration from collapsing into a single soliton. Again, as emphasized above, to the best of our knowledge, our results are the first study of the concentration and fluctuations of Gibbs measures around multi-soliton manifold. In particular, a crucial aspect is to understand how solitons interact with each other on the collision scale $\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, and how the energy behaves at this scale, even though such configurations are not minimizers.
Regarding the phase transition of the discrete focusing NLS Gibbs measure, in [17], for $d\geq 3$, Dey-Kirkpatrick-Krishnan identified a phase transition, analogous to the one found by Tolomeo–Weber [38], on the lattice. Using two parameters, temperature and the strength of the nonlinearity, they proved the existence of a continuous phase transition curve that divides the parameter plane into two regions, the appearance or non-appearance of (single) solitons. In the recent work [26], Krishnan and Ray further investigated the model using the two parameters. They proved that the three regions in the phase diagram lead to three distinct limits. A natural question is whether the discrete (or even continuous) sine–Gordon model, in one dimension or higher, also exhibits a phase transition depending on the temperature and the strength of the coupling constant.
1.3.4. Gibbs measure for stochastic Allen–Cahn equation and single-soliton manifold
For the (invariant) Gibbs measure of the stochastic one-dimensional Allen–Cahn equation, Weber [40] proved that, in a joint low-temperature and infinite-volume limit, the measure concentrates on the single soliton manifold. Subsequently, Otto–Weber–Westdickenberg [34] studied the same limits, with $\varepsilon\to 0$ and $L_{\varepsilon}\to\infty$, but identified the optimal scaling of the interval length $L_{\varepsilon}$ by analyzing the competition between energy $\varepsilon\to 0$ and entropy $L_{\varepsilon}\to\infty$. Under this optimal scaling, they again observed concentration of the measure around the single soliton manifold. Recently, Bertini–Buttà–Di Gesù [6] showed that, beyond the optimal length of the interval $L_{\varepsilon}$, the measure no longer concentrates, and the interfaces (the soliton centers $\xi_{j}$) become asymptotically distributed according to a Poisson point process. It would be interesting to investigate the optimal length scale for which Theorems 1.2, 1.4, and 1.6 remain valid, and to understand what kinds of behavior occur beyond this optimal scale.
1.3.5. Gibbs measure and topological solitons
We refer to the work of Bringmann [8] on topological solitons and Gibbs measures. Bringmann recently studied exterior equivariant wave maps with spatial domain $\mathbb{R}^{3}\setminus B(0,1)$, which admit topological solitons. In contrast to our setting, where the high charge sector $|Q|\geq 2$ has no minimizer, the wave map system admits infinitely many topological sectors, and each sector possesses a unique minimizer. In [8], Bringmann constructed the Gibbs measure on each homotopy class and proved the invariance of this measure under the corresponding PDE flow. Notice that the measure considered by Bringmann [8, (1.10)] takes a form similar to that of the sine–Gordon measure considered here (1.5), (2.1), where the base point is fixed, and the measure is defined on the fluctuation coordinate.
Finally, we remark that, as discussed by Manton–Sutcliffe [29], there exist many models admitting infinitely many disjoint topological sectors, each supporting topological solitons, including the Abelian Higgs, Ginzburg–Landau, and Yang–Mills models. It would be natural to investigate whether the methods developed in this paper can be applied to study the concentration and fluctuations of Gibbs measures in these settings. The present work may be viewed as a first step toward such a program.
2. Notations and preliminaries
2.1. Notations
Throughout the paper, we fix $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$, where $\eta>0$ is an arbitrarily small but fixed number, and we denote by $\langle\cdot,\cdot\rangle$ the $L^{2}$ inner product on the interval $L_{\varepsilon},L_{\varepsilon}]$:
$\langle\phi,\psi\rangle:=\langle\phi,\psi\rangle_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\phi\psi dx,$
where $\phi$ and $\psi$ are real-valued functions. All $L^{p}$-norms appearing below are understood to be over the interval $L_{\varepsilon},L_{\varepsilon}]$, and we suppress the domain from the notation for $1\leq p\leq\infty$
$\|\phi\|_{L^{p}}^{p}=\|\phi\|_{L^{p}([-L_{\varepsilon},L_{\varepsilon}])}^{p}=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\phi|^{p}dx,\quad\|\phi\|_{L^{\infty}}=\|\phi\|_{L^{\infty}([-L_{\varepsilon},L_{\varepsilon}])}.$
When we use $L^{p}$-norms on the real line $\mathbb{R}$, we explicitly write $L^{p}(\mathbb{R})$.
Regarding the topological degree $Q\in\mathbb{Z}$ defined in (1.2), through the paper we mainly consider the case $Q>0$ When $Q<0$, the same results follow by replacing the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}=m_{\xi_{1}}+\dots+m_{\xi_{k}}$, consisting of kinks, with $m_{\xi_{1},\dots,\xi_{k}}^{-}=m_{\xi_{1}}^{-}+\dots+m_{\xi_{k}}^{-}$, consisting of anti-kinks.
Let $A_{1},\dots,A_{k}$ be measurable sets. We use the notation
$\mathbb{E}\Big[F(\phi),A_{1},\dots A_{k}\Big]=\mathbb{E}\bigg[F(\phi)\prod_{j=1}^{k}\mathbf{1}_{A_{j}}\bigg],$
where $\mathbb{E}$ stands for the expectation with respect to the probability distribution of $\phi$ under consideration.
We use $c>0$ to denote an unimportant positive constant whose value may change from line to line. We write $A\lesssim B$ to indicate an estimate of the form $A\leq CB$ for some $C>0$. We also write $A\approx B$ to denote $A\lesssim B$ and $B\lesssim A$ and use $A\ll B$ when we have $A\leq\zeta B$ for some small $\zeta>0$. We may include subscripts to show dependence on external parameters; for example, $A\lesssim_{p}B$ means $A\leq C(p)B$, where the constant $C(p)$ depends on a parameter $p$. In addition, we use $a$- and $a+$ to denote $a-\eta$ and $a+\eta$, respectively for arbitrarily small $\eta>0$.
2.2. Brownian bridges and homotopy-class representatives
Recall that, in defining the base Gaussian measure corresponding to the homotopy class $\mathcal{C}_{Q}$, we introduce the Gaussian measure $\mu^{Q}_{\varepsilon}$ in (1.4), conditioned on $\phi(-L_{\varepsilon})=0$ and $\phi(L_{\varepsilon})=2\pi Q$. This measure is precisely the law induced by the Brownian bridge $B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}$
$B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x),$ (2.1)
where $B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}$ is the mean zero Brownian bridge pinned at 0 at both ends $L_{\varepsilon},L_{\varepsilon}$, and its covariance is given by
$\mathbb{E}\Big[\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{1})\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{2})\Big]=\frac{\varepsilon}{2L_{\varepsilon}}\big((x_{1}+L_{\varepsilon})(L_{\varepsilon}-x_{2})\wedge(x_{2}+L_{\varepsilon})(L_{\varepsilon}-x_{1})\big)$ (2.2)
for all $x_{1},x_{2}\in[-L_{\varepsilon},L_{\varepsilon}]$. In (2.1), $\ell^{Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})$ is the affine function interpolating between 0 and $2\pi Q$. Thus, the fluctuations are of order $\sqrt{\varepsilon}$ around this straight line connecting the boundary values. Recall that we interpret a field $\phi$ distributed according to $\mu^{Q}_{\varepsilon}$ as a function on $\mathbb{R}$, extended by 0 and $2\pi Q$ outside $L_{\varepsilon},L_{\varepsilon}]$. Hence, the reference profile $\ell^{Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})$ is extended by the same boundary values, while the fluctuation coordinate $B^{0,0}_{(-L_{\varepsilon},L_{\varepsilon})}$ is extended by 0 outside $L_{\varepsilon},L_{\varepsilon}]$.
Let $\{e_{n}\}_{n\geq 1}$ be the $L^{2}$-orthonormal eigenfunctions of $\partial_{x}^{2}$ on $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions
$e_{n}(x)=\frac{1}{\sqrt{L_{\varepsilon}}}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big)$ (2.3)
for $n\in\mathbb{N}$. The corresponding eigenvalues are $\lambda_{n}=\Big(\frac{n\pi}{2L_{\varepsilon}}\Big)^{2}$. Then, for $x\in[-L_{\varepsilon},L_{\varepsilon}]$, the Brownian bridge admits the Fourier series representation
$B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}g_{n}}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$ (2.4)
where $\{g_{n}\}_{n\geq 1}$ is a family of independent standard Gaussian random variables.
2.3. Boué-Dupuis formula
In this subsection, we express Gaussian functional integrals with respect to the Brownian bridge measure $\mu^{Q}_{\varepsilon}=\operatorname{Law}\big(B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}\big)$ in (2.1) in terms of an optimal control problem. We first define a centered Gaussian process at each scale $t\in[0,1]$ as follows
$Y(t)=\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)=\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}B_{n}(t)}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$ (2.5)
where $\{B_{n}\}_{n\geq 1}$ is a sequence of independent Brownian motions. Then we have
$\mu^{Q}_{\varepsilon}=\operatorname{Law}\big(B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}\big)=\operatorname{Law}\big((\tfrac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}Y(1)\big).$
Next, let $\mathbb{H}_{a}$ denote the space of drifts, which consists of mean-zero progressively measurable processes belonging to $L^{2}([0,1];L^{2}([-L_{\varepsilon},L_{\varepsilon}]))$, $\mathbb{P}$-almost surely. We are now ready to state the Boué-Dupuis variational formula [7, 39]; in particular, see Theorem 7 in [39]. See also Theorem 2 in [2].
Lemma 2.1.
Let $Q\in\mathbb{Z}$ and let $\varepsilon>0$. Suppose that $F$ is a measurable real-valued functional such that $\mathbb{E}\big[|F(Y(1))|^{p}\big]<\infty$ and $\mathbb{E}\big[|e^{-F(Y(1))}|^{q}\big]<\infty$ for some $1<p,q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$. Then, we have
$-\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\Big[e^{-F(\phi)}\Big]=\inf_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[F\big(\ell^{Q}+\sqrt{\varepsilon}Y(1)+\sqrt{\varepsilon}\Theta(1)\big)+\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg],$
where $\ell^{Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})$ is the line connecting 0 and $2\pi Q$ in (2.1), and
$\Theta(t):=\int_{0}^{t}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta(s)ds.$ (2.6)
Here the expectation $\mathbb{E}=\mathbb{E}_{\mathbb{P}}$ is an expectation with respect to the underlying probability measure $\mathbb{P}$.
In the following, we use the shorthand notations $Y:=Y(1)$ and $\Theta:=\Theta(1)$ for convenience.
In the large deviation estimates (Section 6), we need moment estimates for $Y(t)$ and a pathwise estimate for the drift term.
Lemma 2.2.
Let $Y(t)$ and $\Theta(t)$ be as in (2.5) and (2.6).
(i) For any $t\in[0,1]$, we have
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|^{2}dx\bigg]$ $=\frac{2}{3}t\cdot\varepsilon L_{\varepsilon}^{2},$ (2.7)
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|dx\bigg]$ $=\frac{\sqrt{\pi t}}{2}\cdot\varepsilon^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}.$ (2.8)
(ii) The drift term $\Theta(t)$ has the regularity of the Cameron-Martin space, that is, for any $\theta\in\mathbb{H}_{a}$, we have
$\|\Theta(1)\|_{\dot{H}^{1}_{x}}^{2}\leq\int_{0}^{1}\|\dot{\Theta}(t)\|_{\dot{H}^{1}_{x}}^{2}dt=\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt,$ (2.9)
where $\dot{\Theta}(t)=\langle\nabla\rangle^{-1}\theta(t)$.
Proof.
Part (ii) follows from Minkowski’s and Cauchy-Schwarz’s inequalities. We first prove (2.7). Since $\{e_{n}\}_{n\geq 1}$ is an $L^{2}$-orthonormal basis of $L^{2}([-L_{\varepsilon},L_{\varepsilon}])$, we have $\|e_{n}\|_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}=1$ for all $n\geq 1$. This implies
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|^{2}dx\bigg]$ $=\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\bigg|\sqrt{\varepsilon}\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)\bigg|^{2}dx\bigg]=t\varepsilon\sum_{n\geq 1}\frac{1}{\lambda_{n}}$
$=t\varepsilon(2L_{\varepsilon})^{2}\sum_{n\geq 1}\frac{1}{(n\pi)^{2}}=\frac{2}{3}t\cdot\varepsilon L_{\varepsilon}^{2},$
where we used $\lambda_{n}=(\tfrac{n\pi}{2L_{\varepsilon}})^{2}$. We now prove (2.8). From $Y(t)=\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)$, (2.1), (2.4), and (2.2), we have
$\mathbb{E}|Y(t,x)|^{2}=t\sum_{n\geq 1}\frac{1}{\lambda_{n}}e_{n}(x)^{2}=t\mathbb{E}\Big[B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x)B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x)\Big]=t\frac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}}.$
This implies that for every fixed $t$ and $x$, $\sqrt{\varepsilon}Y(t,x)\sim\mathcal{N}(0,\varepsilon t\tfrac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}})$. Recall that for a centered Gaussian $Z\sim\mathcal{N}(0,\sigma^{2})$, $\mathbb{E}|Z|=\sigma\sqrt{\frac{2}{\pi}}$. Therefore,
$\mathbb{E}|\sqrt{\varepsilon}Y(t,x)|=\sqrt{\frac{2}{\pi}}\sqrt{\varepsilon t\frac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}}}=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\sqrt{L_{\varepsilon}^{2}-x^{2}}.$ (2.10)
Using (2.10), we have
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|dx\bigg]=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sqrt{L_{\varepsilon}^{2}-x^{2}}dx=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\frac{\pi L_{\varepsilon}^{2}}{2}=\frac{\sqrt{\pi}}{2}\varepsilon^{\frac{1}{2}}t^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}.$
3. Structure of multi–topological solitons
In this section, we study the properties of topological solitons and their superpositions forming multi-soliton configurations. The key points are to understand how solitons interact with each other on the collision scale, and how the energy behaves even though these configurations are not minimizers.
3.1. Topological solitons
In this subsection, we investigate the minimizers of the Hamiltonian
$\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
within the homotopy class $\mathcal{C}_{Q}$ for $|Q|=1$, referred to as topological solitons. The topological solitons, namely the kink ($Q=1$) and anti-kink (Q=1Q=-1) centered at $\xi\in\mathbb{R}$
$m_{\xi}(x)=m(x-\xi)=4\arctan(e^{x-\xi}),\qquad m_{\xi}^{-}(x)=m^{-}(x-\xi)=4\arctan(e^{-(x-\xi)})$
are localized transition layers connecting the distinct vacua 0 and $2\pi$, satisfying the Euler–Lagrange equation
$-\partial_{x}^{2}\phi+\sin\phi=0.$
The kink represents an increasing transition from 0 to $2\pi$, whereas the anti-kink corresponds to the decreasing transition from $2\pi$ to 0. These configurations $\{m_{\xi}\}_{\xi\in\mathbb{R}}$ and $\{m_{\xi}^{-}\}_{\xi\in\mathbb{R}}$ minimize the Hamiltonian within their respective topological sectors $\mathcal{C}_{Q}$
$E_{\textup{kink}}=\inf_{\phi\in\mathcal{C}_{1}}E(\phi)=E(m_{\xi})=8,\qquad E_{\textup{anti-kink}}=\inf_{\phi\in\mathcal{C}_{-1}}E(\phi)=E(m_{\xi}^{-})=8.$
In the following lemma, we show that each $m_{\xi}(x)$ is exponentially close to a vacuum (0 or $2\pi$), that is, nearly constant, when $x$ is far from its center $\xi$.
Lemma 3.1.
Let $\xi\in\mathbb{R}$.
  • (i)
    When $x>\xi$,
    $2e^{-|x-\xi|}\leq|m_{\xi}(x)-2\pi|\leq 4e^{-|x-\xi|},\quad 2e^{-|x-\xi|}\leq|m_{\xi}^{-}(x)|\leq 4e^{-|x-\xi|},$
    uniformly in all centers $\xi$ and all $x\in\mathbb{R}$.
  • (ii)
    When $x<\xi$,
    $2e^{-|x-\xi|}\leq|m_{\xi}(x)|\leq 4e^{-|x-\xi|},\quad 2e^{-|x-\xi|}\leq|m_{\xi}^{-}(x)-2\pi|\leq 4e^{-|x-\xi|},$
    uniformly in all centers $\xi$ and all $x\in\mathbb{R}$.
  • (iii)
    We have
    $|\partial_{\xi}m_{\xi}(x)|+|\partial_{\xi}^{2}m_{\xi}(x)|\leq 4e^{-|x-\xi|},\quad|\partial_{\xi}m_{\xi}^{-}(x)|+|\partial_{\xi}^{2}m_{\xi}^{-}(x)|\leq 4e^{-|x-\xi|},$
    uniformly in all centers $\xi$ and all $x\in\mathbb{R}$. The same holds with $\xi$-derivatives replaced by $x$-derivatives.
Proof.
The parts (i) and (ii) follow from direct computation using $m_{\xi}(x)=4\arctan(e^{x-\xi})$ and $m^{-}_{\xi}(x)=4\arctan(e^{-(x-\xi)})$. The part (iii) follows from the fact that $\partial_{\xi}m_{\xi}(x)=-2\textup{sech}(x-\xi)$, and $\partial_{\xi}^{2}m_{\xi}(x)=2\textup{tanh}(x-\xi)\textup{sech}(x-\xi)$ and direct computations. These derivatives are highly localized profile around $\xi$ with an exponentially decaying tail. ∎
In the following lemma, we show that (1) translation is the only symmetry of the minimizer, and (2) if a field $\phi\in\mathcal{C}_{Q}$ is far away from the family of minimizers, then its energy is also far away from the minimal energy.
Lemma 3.2.
Let $|Q|=1$.
  • (1)
    If $G\in\mathcal{C}_{Q}$ satisfies $E(G)=\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)$, then there exists $\xi\in\mathbb{R}$ such that
    $G(x)=m(x-\xi)\quad\textup{when $Q=1$},\quad G(x)=m^{-}(x-\xi)\quad\textup{when $Q=-1$}.$
  • (2)
    Let $\textup{dist}(\phi,\mathcal{M}_{1}):=\inf_{\xi\in\mathbb{R}}\|\phi-m_{\xi}\|_{L^{2}(\mathbb{R})}$ and $\textup{dist}(\phi,\mathcal{M}_{-1}):=\inf_{\xi\in\mathbb{R}}\|\phi-m_{\xi}^{-}\|_{L^{2}(\mathbb{R})}$. There exists $c>0$ such that if $\phi\in\mathcal{C}_{Q}$ satisfies
    $\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta>0,$
    then
    $E(\phi)\geq\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)+c\cdot\textup{dist}(\phi,\mathcal{M}_{Q})^{2}\geq E_{\textup{kink}}+c\cdot\delta^{2}.$
Proof.
The parts (1) and (2) follow from the concentration compactness argument together with
$E(\phi)$ $=E(m_{\xi})+\langle\nabla E(m_{\xi}),\phi-m_{\xi}\rangle_{L^{2}(\mathbb{R})}$
$\qquad+\frac{1}{2}\langle\phi-m_{\xi},\nabla^{2}E(m_{\xi})(\phi-m_{\xi})\rangle_{L^{2}(\mathbb{R})}+O(\|\phi-m_{\xi}\|_{L^{2}(\mathbb{R})}^{3})$
since $\nabla E(m_{\xi})=0$. For details, see, for example, [38, Lemma 2.4], [40, Proposition 2.2], and [20, Lemma 6.5]. ∎
3.2. Multi-topological solitons
Unlike the class $\mathcal{C}_{Q}$ for $|Q|=1$, it is well known that there is no minimizer in the homotopy class $\mathcal{C}_{Q}$ when $|Q|\geq 2$. In this subsection, we therefore investigate the properties of superpositions of topological solitons under appropriate separation conditions, which “almost” act as minimizers in this class (see Remark 3.5). Furthermore, we analyze how the energy behaves when a field is far from the multi-soliton manifold (Lemma 3.7) and when solitons collide on the collision scale (Lemmas 3.10 and 3.12).
First, we study the Bogomolny lower bound on the homotopy class $\mathcal{C}_{Q}$.
Lemma 3.3 (Bogomolny lower bound).
Let $Q\in\mathbb{Z}$. For any $\phi\in\mathcal{C}_{Q}$,
$E(\phi)\geq E_{\textup{kink}}|Q|=8|Q|,$
where $E_{\textup{kink}}=E(m)=E(m^{-})=8$.
Proof.
Note that
$E(\phi)$ $=\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\phi|^{2}dx+\int_{\mathbb{R}}2\sin^{2}\frac{\phi}{2}dx\geq 2\int_{\mathbb{R}}|\partial_{x}\phi\sin\frac{\phi}{2}|dx$
$\geq 2\int_{C}f=\int_{0}^{2\pi|Q|}|\sin(z/2)|dz=8|Q|,$
where $\int_{C}f$ is a line integral, $C$ is the curve in $\mathbb{R}$ going straight from 0 to $2\pi Q$, and $f(z)=|\sin(z/2)|$ is a function along $C$. In the second line we have replaced the parametrization $\phi$ of $C$ by the identity parametrization $0,2\pi Q]\to[0,2\pi Q]$ and used the independence of line integrals on parametrizations. ∎
For $\xi_{1}\leq\cdots\leq\xi_{k}$, where $k=|Q|$, define the superposition of topological solitons, that is, the multi-soliton by
$m_{\xi_{1},\dots,\xi_{k}}$ $=\sum_{j=1}^{k}m(\cdot-\xi_{j})=\sum_{j=1}^{k}m_{\xi_{j}}$
$m_{\xi_{1},\dots,\xi_{k}}^{-}$ $=\sum_{j=1}^{k}m^{-}(\cdot-\xi_{j})=\sum_{j=1}^{k}m^{-}_{\xi_{j}}.$
In the following lemmas, we state the results only for the case $Q>0$, by symmetry.
We first prove an elementary inequality: for $\eta\in(0,1)$,
$\int_{\mathbb{R}}e^{-(|y-x|+|z-x|)}dx=(1+|y-z|)e^{-|y-z|}\leq\frac{1}{\eta}e^{\eta-1}e^{-(1-\eta)|y-z|}.$ (3.1)
The equality follows by observing that the integral between $y$ and $z$ is equal to $|y-z|e^{-|y-z|}$, and the integral outside is equal to $e^{-|y-z|}$. For the inequality, we have
$(1+|y-z|)e^{-|y-z|}\leq\Big(\sup_{r>0}(1+r)e^{-\eta r}\Big)e^{-(1-\eta)|y-z|}$
and the function $(1+r)e^{-\eta r}$ reaches maximum at $r=\frac{1}{\eta}-1$.
When the centers $\xi_{1},\dots,\xi_{k}$ are well separated, each kink $m_{\xi_{j}}$ contributes its own $E_{\textup{kink}}$, and the overlap between kinks yields only exponentially small corrections, because each kink remains nearly constant (0 or $2\pi$) outside its center $\xi_{j}$. This is shown in the next lemma.
Lemma 3.4.
Let $Q\in\mathbb{Z}$ with $|Q|=k$. Then we have
$E(m_{\xi_{1},\dots,\xi_{k}})$ $=\sum_{j=1}^{k}E(m_{\xi_{j}})+O(e^{-c\min_{i\neq j}|\xi_{i}-\xi_{j}|})$
as $\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$, where $c>0$ only depends on $|Q|$.
Proof.
The kinetic energy part is
$\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}m_{\xi_{1},\dots,\xi_{k}}|^{2}dx$ $=\frac{1}{2}\sum_{j=1}^{k}\int_{\mathbb{R}}|\partial_{x}m_{\xi_{j}}|^{2}dx+\sum_{i<j}\int_{\mathbb{R}}\partial_{x}m_{\xi_{i}}\partial_{x}m_{\xi_{j}}dx$
$=\frac{1}{2}\sum_{j=1}^{k}\int_{\mathbb{R}}|\partial_{x}m_{\xi_{j}}|^{2}dx+\sum_{i<j}O(e^{-c|\xi_{i}-\xi_{j}|})$
by Lemma 3.1(iii) and (3.1).
We now study the potential energy part. Let $U(z)=1-\cos(z)$. One has
$|U(a+b)-U(a)-U(b)|=\Big|\int_{0}^{a}\int_{0}^{b}U^{\prime\prime}(s+t)dsdt\Big|\leq|a||b|$
and by induction in $k$ one has
$\Big|U\Big(\sum_{j=1}^{k}a_{j}\Big)-\sum_{j=1}^{k}U(a_{j})\Big|\leq\sum_{i<j}|a_{i}||a_{j}|.$ (3.2)
Indeed, suppose that this holds for k1k-1. Then the left-hand side of (3.2) is bounded by
$\Big|U\Big(\sum_{j=1}^{k}a_{j}\Big)-U\Big(\sum_{j=1}^{k-1}a_{j}\Big)-U(a_{k})\Big|+\Big|U\Big(\sum_{j=1}^{k-1}a_{j}\Big)+U(a_{k})-\sum_{j=1}^{k}U(a_{j})\Big|$
and (3.2) follows by using the induction assumption.
Set $a_{j}=m_{\xi_{j}}(x)-2\pi 1_{x>\xi_{j}}$. (Note that the shift $2\pi 1_{x>\xi_{j}}$ does not change the value of $\cos$.) By Lemma 3.1(i)(ii), $|a_{j}|\leq 4e^{-|x-\xi_{j}|}$. The desired bound follows upon integrating over $x$ again by (3.1). ∎
Remark 3.5.
According to the Bogomolny structure in Lemmas 3.3 and 3.4, by choosing a minimizing sequence with infinite separation between centers, we can show that the minimal energy in each homotopy class $\mathcal{C}_{Q}$ is
$\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)=|Q|E_{\textup{kink}}.$
This shows that the infimum of the energy in the topological sector $\mathcal{C}_{Q}$ is attained asymptotically by configurations consisting of $Q$ kinks when $Q>0$ or $|Q|$ anti-kinks when $Q<0$, with infinite mutual separation $\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$. However, the minimizing sequence has no convergent subsequence due to the infinite separation, which cannot be realized in practice. This is one of the reasons why there is no actual minimizer for $|Q|\geq 2$.
From Remark 3.5, under the separation condition $m_{\xi_{1},\dots,\xi_{k}}$ is an almost minimizer on the homotopy class $\mathcal{C}_{Q}$ (though it never attains the actual minimum). In the following lemma, we show that the first variation operator is close to zero.
Lemma 3.6.
Let $Q\in\mathbb{Z}$ with $|Q|=k$. There exists $0<\eta\ll 1$ sufficiently small such that
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq C^{\nabla}\frac{e^{-(1-\eta)d}}{1-e^{-(1-\eta)d}}$
uniformly in $\xi_{1},\dots,\xi_{k}$ where $d:=\min_{i\neq j}|\xi_{i}-\xi_{j}|$ and $C^{\nabla}=(\frac{1}{2\eta}e^{\eta-1})^{1/2}Q$.
Proof.
Note that
$\nabla E(m_{\xi_{1},\dots,\xi_{k}})(x)=-\partial_{x}^{2}m_{\xi_{1},\dots,\xi_{k}}(x)+\sin m_{\xi_{1},\dots,\xi_{k}}(x)=\sin m_{\xi_{1},\dots,\xi_{k}}(x)-\sum_{j=1}^{k}\sin m_{\xi_{j}}(x)$
where we used $\nabla E(m_{\xi_{j}})=-\partial_{x}^{2}m_{\xi_{j}}+\sin m_{\xi_{j}}=0$. Note that (3.2) again holds with $U(z)=\sin(z)$. So
$|\nabla E(m_{\xi_{1},\dots,\xi_{k}})(x)|\leq\sum_{i<j}e^{-|x-\xi_{i}|}e^{-|x-\xi_{j}|}.$ (3.3)
Regarding the $L^{2}$ bound, note that by (3.1),
$\int_{\mathbb{R}}e^{-2\,(|x-\xi_{i}|+|x-\xi_{j}|)}dx\leq\frac{1}{2\eta}e^{\eta-1}e^{-2(1-\eta)|\xi_{i}-\xi_{j}|}$
for small $\eta>0$. This implies
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}\sum_{i<j}e^{-(1-\eta)|\xi_{i}-\xi_{j}|}.$
One could estimate $\sum_{i<j}$ by $Q^{2}$, but to be slightly sharper, we assume without loss of generality that $\xi_{1}\leq\cdots\leq\xi_{k}$. Then there are Q1Q-1 terms of the form $e^{-(1-\eta)|\xi_{i}-\xi_{i+1}|}$, and Q2Q-2 terms of the form $e^{-(1-\eta)|\xi_{i}-\xi_{i+2}|}$ which can be written as $e^{-(1-\eta)|\xi_{i}-\xi_{i+1}|-(1-\eta)|\xi_{i+1}-\xi_{i+2}|}$, etc. Thus
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}\sum_{k=1}^{Q-1}(Q-k)e^{-(1-\eta)kd}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}Q\frac{e^{-(1-\eta)d}}{1-e^{-(1-\eta)d}}.$
The following lemma shows that if a field $\phi\in\mathcal{C}_{Q}$ is far away from the multi-soliton manifold, then its energy is also far away from the minimal energy.
Lemma 3.7.
Let $Q\in\mathbb{Z}$ with $Q\geq 2$ and let $\delta>0$. Then there exists $c>0$, depending only on $Q$, such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
$\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}\geq\delta>0,$ (3.4)
then
$E(\phi)\geq\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)+c\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}\geq\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)+c\delta^{2}.$
Proof.
Fix $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$. Define
$t_{0}=-\infty,\quad t_{Q}=\infty,\quad t_{j}=\inf\{x:\phi(x)=2\pi j\}$
for $1\leq j\leq Q-1$. Then, define on each block $1\leq j\leq Q$
$\phi_{j}(x)=\begin{cases}\phi(x)\quad&x\in(t_{j-1},t_{j})\\ 2\pi(j-1)\quad&x<t_{j-1}\\ 2\pi j\quad&x>t_{j}.\end{cases}$ (3.5)
Since $2\pi\mathbb{Z}$ extension does not change the energy, we have, for $1\leq j\leq Q$,
$E(\phi_{j})=\int_{t_{j-1}}^{t_{j}}|\partial_{x}\phi|^{2}dx+\int_{t_{j-1}}^{t_{j}}(1-\cos\phi)dx.$ (3.6)
Therefore, using (3.6) and Lemma 3.2-(2), we have
$E(\phi)=\sum_{j=1}^{Q}E(\phi_{j})$ $=\sum_{j=1}^{Q}E(\phi_{j}-2\pi(j-1))$
$\geq|Q|E_{\textup{kink}}+c\sum_{j=1}^{Q}\inf_{\xi\in\mathbb{R}}\|(\phi_{j}-2\pi(j-1))-m_{\xi}\|_{L^{2}(\mathbb{R})}^{2}.$ (3.7)
In the following, we show that
$\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}\leq c\sum_{j=1}^{Q}d_{j}^{2},$ (3.8)
where $d_{j}:=\inf_{\xi\in\mathbb{R}}\|(\phi_{j}-2\pi(j-1))-m_{\xi}\|_{L^{2}(\mathbb{R})}$. By the definition of $d_{j}$, for each $j$ we can pick $\xi_{j}^{*}\in\mathbb{R}$ so that
$\|(\phi_{j}-2\pi(j-1))-m_{\xi_{j}^{*}}\|_{L^{2}(\mathbb{R})}^{2}<d_{j}^{2}+\varepsilon$ (3.9)
for some small $\varepsilon>0$ (later we let $\varepsilon\to 0$). Fix such a choice $(\xi_{1}^{*},\dots,\xi_{Q}^{*})$ from now on. By the definition of $\phi_{j}$, one can easily check that for all $x\in\mathbb{R}$
$\phi(x)=\sum_{j=1}^{Q}(\phi_{j}(x)-2\pi(j-1)).$
This, together with (3.9), implies that
$\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}$ $\leq 2^{Q-1}\sum_{j=1}^{Q}\|(\phi_{j}-2\pi(j-1))-m_{\xi_{j}^{*}}\|_{L^{2}(\mathbb{R})}^{2}$
$\leq 2^{Q-1}\sum_{j=1}^{Q}d_{j}^{2}+2^{Q-1}|Q|\varepsilon.$
Letting $\varepsilon\to 0$, we obtain (3.8). Combining (3.7) and (3.8) yields
$E(\phi)\geq|Q|E_{\textup{kink}}+c\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}.$
Since $\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)=|Q|E_{\textup{kink}}$ from Remark 3.5, we obtain the desired result.
Remark 3.8.
In Lemma 3.7, we obtain an energy gap estimate when the field is far from the multi-soliton manifold $\mathcal{M}^{Q}$. The main point is to quantify how much the energy exceeds the minimal energy. Specifically, the energy is higher than the minimum by an amount of order $\delta^{2}$ when $\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta$.
Remark 3.9.
In Lemma 3.7, when $Q\in\mathbb{Z}$ is negative with $Q\leq-2$, the same result holds by replacing the condition (3.4) with the multi-antikink manifold, as follows:
$\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{|Q|}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}^{-}\|_{L^{2}(\mathbb{R})}\geq\delta>0,$
where $m_{\xi_{1},\dots,\xi_{k}}^{-}=m^{-}(\cdot-\xi_{1})+\dots+m^{-}(\cdot-\xi_{|Q|})$.
Recall from Remark 3.5 that the minimal energy $\inf_{\phi\in\mathcal{C}_{Q}}E(\phi)$ is achieved in the limit of multi-solitons with infinite separation, that is, when $\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$. In the following lemma, we show that although a field $\phi$ may be close to the multi-soliton manifold $\textup{dist}(\phi,\mathcal{M}^{Q})<\delta$, if the solitons are not well separated $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, then its energy remains far from the minimal energy.
Before beginning the proof, we recall that $\mathcal{M}^{<d}_{Q}$ is the collision manifold introduced in (1.9). The main part is to quantify the error, namely, how much the energy exceeds the minimal energy $|Q|E_{\textup{kink}}$ when the solitons collide on the scale $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$. In this regime, the energy is higher than the minimum by an amount of order $e^{-d}$.
Lemma 3.10.
Let $Q\in\mathbb{Z}$ with $Q\geq 2$, and let $d>0$ be a large constant. Then there exists $c>0$ such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
$\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})=\inf_{\begin{subarray}{c}\xi_{1},\dots,\xi_{Q}\in\mathbb{R}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|<d\end{subarray}}\|\phi-m_{\xi_{1},\dots,\xi_{Q}}\|_{L^{2}(\mathbb{R})}<\delta$ (3.10)
where $\frac{1}{2}e^{-d}\geq c\cdot\delta$, then
$E(\phi)\geq|Q|E_{\textup{kink}}+c\cdot e^{-d}.$
Proof.
Since $\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta$, we can find $\xi_{1}^{*},\dots,\xi_{Q}^{*}$ such that
$\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|<d\quad\textup{and}\quad\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta.$ (3.11)
Therefore, from Lemma 3.12, we obtain
$E(\phi)$ $=\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$\geq\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+|Q|E_{\textup{kink}}+c\cdot e^{-d}.$ (3.12)
By denoting $\eta(x):=\phi(x)-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)$, we write
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $=\int\partial_{x}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}(1-\cos(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\eta))dx-\int_{\mathbb{R}}(1-\cos m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})dx.$ (3.13)
Use Taylor on the cosine pointwise
$\cos(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)+\eta(x))$
$=\cos m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)-\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)\cdot\eta(x)-\frac{1}{2}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)+\theta(x)\eta(x)\big)\cdot\eta(x)^{2}$ (3.14)
for some $\theta(x)\in(0,1)$. By plugging (3.14) into (3.13), we obtain
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $=\int\partial_{x}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\cdot\eta dx+\frac{1}{2}\int_{\mathbb{R}}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\theta\eta\big)\cdot\eta^{2}dx.$
Integrating by parts,
$E(\phi)$ $-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$=\int_{\mathbb{R}}\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\cdot\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx+\frac{1}{2}\int_{\mathbb{R}}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\theta\eta\big)\cdot\eta^{2}dx,$
where $\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})=-\partial_{x}^{2}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}$. Using $\|\eta\|_{L^{2}(\mathbb{R})}=\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta$ and $\cos\geq-1$,
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\|\eta\|_{L^{2}(\mathbb{R})}^{2}$
$\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\delta^{2}.$ (3.15)
From (3.3), we have $\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\lesssim e^{-(1-\eta)\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|}$. This implies that
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\geq-c\delta-\frac{1}{2}\delta^{2}\geq-c_{1}\delta$ (3.16)
for all $0<\delta\leq\delta_{0}$, where $\delta_{0}>0$ is sufficiently small and $c_{1}>0$. Combining (3.16) and (3.12) yields
$E(\phi)\geq|Q|E_{\textup{kink}}+c\cdot e^{-d}-c_{1}\delta\geq|Q|E_{\textup{kink}}+\frac{c}{2}\cdot e^{-d}$ (3.17)
under the condition $\frac{c}{2}e^{-d}\geq c_{1}\delta$.
Remark 3.11.
In Lemma 3.10, when $Q\in\mathbb{Z}$ is negative with $Q\leq-2$, the same result holds by replacing the condition (3.10) with the multi-antikink manifold.
It remains to prove Lemma 3.12, which was used in the proof of Lemma 3.10. As alluded above, the key is to quantify the error, namely, how much the energy exceeds the minimal energy in the collision regime. Unlike the previous results such as Lemma 3.4 and Lemma 3.6 where kink interactions are weak when well-separated, here in the collision regime their interactions are stronger. In order to better understand their interactions in this regime, our analysis will have a more “pointwise” (rather than $L^{2}$) flavor, and we will prove an exact pointwise representation for the energy density gap. From this representation we will be able to identify the leading order contribution to the energy gap.
Lemma 3.12.
Let $Q\in\mathbb{Z}$ with $Q\geq 2$. For $d>0$ sufficiently large, there exists $c>0$ such that for all $\xi_{1},\dots,\xi_{Q}\in\mathbb{R}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, one has
$E(m_{\xi_{1},\dots,\xi_{Q}})\geq|Q|E_{\textup{kink}}+c\cdot e^{-d}.$
Proof.
From Lemma 3.1, we have two-sided exponential asymptotic tail: for $x<\xi$,
$2e^{-|x-\xi|}\leq m_{\xi}(x)=4\arctan(e^{x-\xi})\leq 4e^{-|x-\xi|}.$ (3.18)
Similar lower bound holds for $x>\xi$. Recall from the proof of Lemma 3.4 that
$E$ $(m_{\xi_{1},\dots,\xi_{Q}})-|Q|E_{\textup{kink}}$
$=\int_{\mathbb{R}}\sum_{i<j}\partial_{x}m_{\xi_{i}}\partial_{x}m_{\xi_{j}}+1-Q-\cos\Big(\sum_{j=1}^{Q}m_{\xi_{j}}\Big)+\sum_{j=1}^{Q}\cos(m_{\xi_{j}})\;dx.$
It is elementary to check that (let $\tan\theta=e^{x-\xi_{i}}$ and use $\sin(2\theta)=\frac{2\tan\theta}{1+(\tan\theta)^{2}}$)
$\partial_{x}m_{\xi_{i}}=\frac{4e^{x-\xi_{i}}}{1+e^{2(x-\xi_{i})}}=2\sin(2\arctan(e^{x-\xi_{i}}))=2\sin(m_{\xi_{i}}(x)/2).$
We claim that
$F_{Q}$ $(\xi_{1},\cdots,\xi_{Q},x)$
$:=4\sum_{i<j}\sin(m_{\xi_{i}}(x)/2)\sin(m_{\xi_{j}}(x)/2)+1-Q-\cos\Big(\sum_{j=1}^{Q}m_{\xi_{j}}(x)\Big)+\sum_{j=1}^{Q}\cos(m_{\xi_{j}}(x))\geq 0$
for all $x\in\mathbb{R}$ and all $\xi_{1}\leq\cdots\leq\xi_{Q}$. In other words the energy density of $m_{\xi_{1},\dots,\xi_{Q}}$ is pointwisely greater than the sum of the energy densities of the corresponding single kinks.
To prove the claim, we first note that $F_{1}=0$. Suppose we have proved that $F_{Q-1}(\xi_{1},\cdots,\xi_{Q-1},x)\geq 0$ for all $x\in\mathbb{R}$ and all $\xi_{1}\leq\cdots\leq\xi_{Q-1}$, and we now prove this for $F_{Q}$. Consider the case $x\leq\xi_{Q}$ (the other possibility is $x\geq\xi_{1}$ which is similar). It suffices to prove
$F_{Q}(\xi_{1},\cdots,\xi_{Q},x)\geq F_{Q}(\xi_{1},\cdots,\xi_{Q-1},\infty,x)$
since the right-hand side is equal to $F_{Q-1}(\xi_{1},\cdots,\xi_{Q-1},x)$, noting that $m_{\infty}(x)=0$. To this end, writing $m_{j}=m_{\xi_{j}}(x)/2$, one has
$F_{Q}$ $(\xi_{1},\cdots,\xi_{Q},x)-F_{Q}(\xi_{1},\cdots,\xi_{Q-1},\infty,x)$
$=\Big(4\sin m_{Q}\sum_{j=1}^{Q-1}\sin m_{j}-\cos(2m_{Q}+2A)+\cos(2m_{Q})\Big)-\Big(\cos 2A-1\Big)$
$=4\sin m_{Q}\sum_{j=1}^{Q-1}\sin m_{j}-2\sin^{2}m_{Q}+2\sin(m_{Q}+2A)\sin(m_{Q})$
$=4\sin m_{Q}\Big(\sum_{j=1}^{Q-1}\sin m_{j}+\cos(m_{Q}+A)\sin A\Big)$
where $A:=\sum_{j=1}^{Q-1}m_{j}$ and we have used sum-to-product trigonometric identities. Note that $\sum_{j=1}^{Q-1}\sin m_{j}\geq|\sin(\sum_{j=1}^{Q-1}m_{j})|$ for any $m_{1},\cdots,m_{Q-1}\in[0,\pi]$, which holds by the induction. Since $\cos(m_{Q}+A)\in[-1,1]$, the above expression is indeed non-negative. Thus we have proved $F_{Q}(\xi_{1},\cdots,\xi_{Q},x)\geq 0$.
Figure 2. The solid curve plots $F_{3}(\xi_{1},\xi_{2},\xi_{3},\cdot)$ with $(\xi_{1},\xi_{2},\xi_{3})=(3,5,8)$. Clearly, the main contribution to the energy gap is in the neighborhoods of $\xi_{j}$. The dashed curve plots $F_{3}(\xi_{1},\xi_{2},\infty,\cdot)$.
Now we need to prove that $\int_{\mathbb{R}}F_{Q}(\xi_{1},\cdots,\xi_{Q},x)dx\geq ce^{-\dot{d}}$. To this end we show the following representation of $F_{Q}$
$F_{Q}(\xi_{1},\cdots,\xi_{Q},x)$
$:=4\sum_{i<j}\sin m_{i}\sin m_{j}+1-Q-\cos\Big(2\sum_{j=1}^{Q}m_{j}\Big)+\sum_{j=1}^{Q}\cos(2m_{j})$
$=4\sum_{i<j}\sin m_{i}\sin m_{j}(\cos(m_{i}+m_{j})+1)$ (3.19)
$\qquad+\sum_{n=3}^{Q}2^{n}\sum_{i_{1}<\cdots<i_{n}}\sin m_{i_{1}}\cdots\sin m_{i_{n}}$
$\qquad\qquad\times\Big((-1)^{\frac{n-1}{2}}\sin(m_{i_{1}}+\cdots+m_{i_{n}})1_{n\in 2\mathbb{Z}-1}+(-1)^{\frac{n-2}{2}}\cos(m_{i_{1}}+\cdots+m_{i_{n}})1_{n\in 2\mathbb{Z}}\Big).$
This representation may be of independent interest; [^7] It is somehow reminiscent with Brydges–Kennedy expansion [10] for 2D sine-Gordon. however below we will mainly use the fact that the “2-body interaction” terms (which have good positive sign since $m_{i}\in[0,\pi]$) exhibit the desired lower bound whereas the $\geq 3$-body interaction” terms (which do not have good sign) will have much smaller absolute values since they have more $\sin$ factors, if we observe in the right regime.
Assume that the above identity holds. Suppose that $|\xi_{\ell}-\xi_{\ell+1}|=d$ is the smallest distance. Thanks to the above pointwise positivity, it suffices to prove the lower bound for the integral of $F_{Q}$ over $x\in[\xi_{\ell}-1,\xi_{\ell}+1]$. Fix $x$ in this interval.
By the lower bound in (3.18), $\sin m_{\ell}=\sin(m_{\xi_{\ell}}(x)/2)\geq\sin(e^{-1})\geq\frac{1}{2e}$. Also, for every $j\neq\ell$, again by the lower bound in (3.18),
$\sin m_{j}=\sin(m_{\xi_{j}}(x)/2)\geq\sin(e^{-|x-\xi_{\ell}|})\geq\frac{1}{2}e^{-|\xi_{j}-\xi_{\ell}|}$
and since $|\xi_{j}-x|>d-1$, $m_{\xi_{j}}(x)$ is close to 0 or $2\pi$ up to an error bounded by $4e^{-(d-1)}$ using the upper bound in (3.18). Therefore,
$\sin m_{j}\leq 2e^{-(d-1)}.$
On the other hand, since $x$ is in a neighborhood around $\xi_{\ell}$, $m_{\xi_{\ell}}(x)/2$ is in a neighborhood of $\pi/2$. More precisely, there is a universal constant $c_{0}>0$ (as long as $d>10$) such that
$\cos(m_{\ell}+m_{j})+1\geq c_{0}$
since $m_{j}$ is close to 0 or $\pi$ up to an exponentially small error as shown above.
Summarizing these bounds, the “2-body interaction” terms
$4\sum_{i<j}\sin m_{i}\sin m_{j}(\cos(m_{i}+m_{j})+1)$ $\geq 4\sin m_{\ell}\sin m_{\ell+1}(\cos(m_{\ell}+m_{\ell+1})+1)$
$\geq c_{2}e^{-|\xi_{\ell+1}-\xi_{\ell}|}=c_{2}e^{-d}$
for $c_{2}=4\cdot(2e)^{-1}\cdot\frac{1}{2}\cdot c_{0}>0$. For the “$n$-body interaction” terms with $n\geq 3$, using the above bounds on $\sin m_{j}$,
$\Big|\sin m_{i_{1}}\cdots\sin m_{i_{n}}\Big|\leq(2e^{-(d-1)})^{n-1},$
which, even with $\sum_{n=3}^{Q}2^{n}\sum_{i_{1}<\cdots<i_{n}}$, is much smaller than $c_{2}e^{-d}$ for $d>0$ sufficiently large.
So the proof is complete once we verify the identity (3.19). We will use the following product-to-sum identities: if $n$ is even,
$\prod_{k=1}^{n}\sin\theta_{k}=\frac{(-1)^{n/2}}{2^{n}}\sum_{e}\cos(e_{1}\theta_{1}+\cdots+e_{n}\theta_{n})\prod_{j=1}^{n}e_{j}$
and if $n$ is odd,
$\prod_{k=1}^{n}\sin\theta_{k}=\frac{(-1)^{(n-1)/2}}{2^{n}}\sum_{e}\sin(e_{1}\theta_{1}+\cdots+e_{n}\theta_{n})\prod_{j=1}^{n}e_{j}$
where $e$ sums over $e=(e_{1},\cdots,e_{n})\in\{1,-1\}^{n}$.
On the RHS of (3.19), by the above identity, for $n$ odd,
$($- $1)^{\frac{n-1}{2}}2^{n}\sin m_{i_{1}}\cdots\sin m_{i_{n}}\sin(m_{i_{1}}+\cdots+m_{i_{n}})$
$=\sum_{e}\sin(e_{1}m_{i_{1}}+\cdots+e_{n}m_{i_{n}})\sin(m_{i_{1}}+\cdots+m_{i_{n}})\prod_{j=1}^{n}e_{j}$
$=\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})-\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
$=-\sum_{e}\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$ (3.20)
where the last step is by a change of variables $e_{j}\to-e_{j}$, and similarly for $n$ even,
$($- $1)^{\frac{n-2}{2}}2^{n}\sin m_{i_{1}}\cdots\sin m_{i_{n}}\cos(m_{i_{1}}+\cdots+m_{i_{n}})$
$=-\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})+\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
which actually gives the same result (3.20) by a change of variables.
For $n=Q$, the $e=(1,\cdots,1)$ term precisely gives us the term $\cos(2\sum_{j=1}^{Q}m_{j})$ on the LHS of (3.19). Consider without loss of generality $\cos(2m_{1}+\cdots+2m_{k})$ for $2\leq k<Q$. This term shows up once in the case $n=Q$ and $e=(1,\cdots,1,-1,\cdots,-1)$ with $\prod_{j}e_{j}=(-1)^{Q-k}$, and shows up $Q-k\choose 1$} times in the case n=Q1n=Q-1 and $e=(1,\cdots,1,-1,\cdots,-1)$ with $\prod_{j}e_{j}=(-1)^{Q-k-1}$, etc., and finally it shows up once in the case $n=k$ and $e=(1,\cdots,1)$. So by binomial identity $(1-1)^{Q-k}=0$ which means there is no such term $\cos(2m_{1}+\cdots+2m_{k})$ appearing on the LHS. Regarding the case $k=1$, the same argument applies except that we do not have the “final” case $n=k=1$ because we only sum over $n\geq 2$: this precisely gives us the term $\cos(2m_{1})$ on the LHS. Then for each $n$ we also have a term with $e_{j}=-1$ for all $j$, and they precisely sum up to the constant 1Q1-Q on the LHS of (3.19). ∎
Before concluding this subsection, we present the following quantified energy gap estimate for the middle-range case.
Lemma 3.13.
Let $Q\in\mathbb{Z}$ with $Q\geq 2$, and let $d>0$ be a large constant. Then there exists $c>0$ such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
$\textup{dist}(\phi,\mathcal{M}_{Q}^{d\leq\cdot<(1+\eta)d}):=\inf_{\begin{subarray}{c}\xi_{1},\dots,\xi_{Q}\in\mathbb{R}\\ d\leq\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d\end{subarray}}\|\phi-m_{\xi_{1},\dots,\xi_{Q}}\|_{L^{2}(\mathbb{R})}<\delta$ (3.21)
where $\frac{1}{2}e^{-d}\geq c\cdot\delta$, then
$E(\phi)\geq|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}.$
Proof.
The proof is almost the same as that of Lemma 3.10. We only present the different parts. By following (3.11) and (3.15), we can find $\xi_{1}^{*},\dots,\xi_{Q}^{*}$ such that
$d\leq\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|<(1+\eta)d\quad\textup{and}\quad\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta$
and
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\delta^{2}.$
Compared to (3.16), we have control of $\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}$ thanks to the condition $d\leq\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|$, which implies
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-e^{-(1-\eta)d}\cdot\delta-\frac{1}{2}\delta^{2},$
where we used Lemma 3.6. Then, under the condition $\frac{1}{2}e^{-d}\geq c\cdot\delta$ and Lemma 3.12,
$E(\phi)$ $=\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$\geq\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}$
$\geq|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}-e^{-(1-\eta)d}\cdot\delta-\frac{1}{2}\delta^{2}$
$\geq|Q|E_{\textup{kink}}+\frac{c}{2}\cdot e^{-(1+\eta)d}.$
This completes the proof.
4. Geometry of the multi-soliton manifold
4.1. Approximate multi-soliton manifold
In the previous section, we studied the multi-solitons defined on $\mathbb{R}$ and their properties. When restricted to the interval $L_{\varepsilon},L_{\varepsilon}]$, where $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}\to\infty$ as $\varepsilon\to 0$, we define an approximating soliton $m^{\varepsilon}_{\xi}$ and multi-solitons $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$, which becomes an increasingly accurate approximation of the soliton $m_{\xi}$ and the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}$ on $\mathbb{R}$ as $\varepsilon\to 0$.
Define $m^{\varepsilon}$ to be a smooth, monotone function that coincides with $m$ on $\varepsilon^{-\frac{1}{2}+2\eta},\varepsilon^{-\frac{1}{2}+2\eta}]$, where $\varepsilon^{-\frac{1}{2}+2\eta}\ll L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$, and is extended to the constants 0 and $2\pi$ outside a slightly larger interval:
$m^{\varepsilon}(x)=\begin{cases}m(x),\;&x\in[-\varepsilon^{-\frac{1}{2}+2\eta},\varepsilon^{-\frac{1}{2}+2\eta}],\\ 2\pi,\;&x\geq\varepsilon^{-\frac{1}{2}+2\eta}+1,\\ 0,\;&x\leq-\varepsilon^{-\frac{1}{2}+2\eta}-1.\end{cases}$
On the transition intervals $\varepsilon^{-\frac{1}{2}+2\eta},\varepsilon^{-\frac{1}{2}+2\eta}+1]$ and $\varepsilon^{-\frac{1}{2}+2\eta}-1,-\varepsilon^{-\frac{1}{2}+2\eta}]$, we require
$m(x)\leq m^{\varepsilon}(x)\leq 2\pi\quad\textup{and}\quad m(x)\geq m^{\varepsilon}(x)\geq 0.$
Define the translated soliton
$m^{\varepsilon}_{\xi}(x):=m^{\varepsilon}(x-\xi),$ (4.1)
for $\xi\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$, where
$\overline{L}_{\varepsilon}:=L_{\varepsilon}-\varepsilon^{-\frac{1}{2}+2\eta}-1=L_{\varepsilon}(1-\varepsilon^{\eta}-\varepsilon^{\frac{1}{2}-\eta})\sim L_{\varepsilon}.$ (4.2)
This choice of $\overline{L}_{\varepsilon}$ ensures that the entire transition region of the kink $m_{\xi}^{\varepsilon}$ (from 0 to $2\pi$) remains inside the interval $L_{\varepsilon},L_{\varepsilon}]$. Note that one can easily check that, for any $1\leq p\leq\infty$
$\|m_{\xi}-m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (4.3)
$\|\partial_{x}m_{\xi}-\partial_{x}m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}.$ (4.4)
Therefore, as $\varepsilon\to 0$, $m_{\xi}^{\varepsilon}$ becomes a more and more precise approximation of the topological soliton $m_{\xi}$ on $\mathbb{R}$. We now define the multi-soliton, given by the superposition of single solitons
$m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m^{\varepsilon}_{\xi_{j}},$ (4.5)
where $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. When $k=Q$ with $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$, this definition is enough for the multi-soliton $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ to do the transition from 0 to $2\pi Q$ on the interval $L_{\varepsilon},L_{\varepsilon}]$. From the single-soliton bounds (4.3) and (4.4), we immediately get the same type of exponential approximation for the multi-soliton
$\|m_{\xi_{1},\dots,\xi_{k}}-m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (4.6)
$\|\partial_{x}m_{\xi_{1},\dots,\xi_{k}}-\partial_{x}m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}},$ (4.7)
where $1\leq p\leq\infty$ and $\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. Thanks to (4.6) and (4.7), we can transfer all lemmas in Section 3 from the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}$ to the approximating multi-soliton $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ as $\varepsilon\to 0$.
With the definition of the multi-soliton in (4.5), we define the (approximate) multi-soliton manifold as follows
$\mathcal{M}_{k}^{\varepsilon}:=\big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{k}\leq\overline{L}_{\varepsilon}\big\},$ (4.8)
where $\overline{L}_{\varepsilon}$ is defined as in (4.2). Then $\mathcal{M}^{\varepsilon}_{k}$ is a manifold of dimension $k$. Inside the multi-soliton manifold $\mathcal{M}^{\varepsilon}_{k}$, we further decompose into the collision region
$\mathcal{M}^{\varepsilon,<d}_{k}:=\Big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\},$ (4.9)
where the solitons interact with each other at distances less than $d$, and the non-collision region
$\mathcal{M}^{\varepsilon,\geq d}_{k}:$ $=\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\},$ (4.10)
where the interaction between solitons is negligible as $d\to\infty$. Note that $\mathcal{M}^{\varepsilon,\geq d}_{k}$ is a closed subset of the compact set $\mathcal{M}^{\varepsilon}_{k}$, and therefore $\mathcal{M}^{\varepsilon,\geq d}_{k}$ is also compact.
4.2. Tangent and normal spaces of the manifold and disintegration formula
In this subsection, we study the geometry of the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq d}_{k}$, its tangent and normal spaces, and a related disintegration formula.
Remark 4.1.
Notice that the multi-soliton manifold $\mathcal{M}_{k}^{\varepsilon}$ and the collision manifold $\mathcal{M}_{k}^{\varepsilon,<d}$, defined in (4.8) and (4.9), are topological manifolds but fail to be differentiable manifolds for $k\geq 3$ due to the presence of collision strata.
For example, when $k=3$, the parameter space $\{\xi_{1}\leq\xi_{2}\leq\xi_{3}\}$ has two codimension-one faces, given by $\xi_{1}=\xi_{2}$ and $\xi_{2}=\xi_{3}$. Each point on these faces admits a neighborhood that is locally diffeomorphic to $\mathbb{R}_{+}\times\mathbb{R}^{2}$. However, there is also an edge corresponding to $\xi_{1}=\xi_{2}=\xi_{3}$ and each point on this edge has a neighborhood that is locally diffeomorphic to $\mathbb{R}^{2}_{+}\times\mathbb{R}$. Since $\mathbb{R}^{2}_{+}$ is the quadrant with a corner, which is homeomorphic but not diffeomorphic to the half-plane, $\mathcal{M}_{k}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ fail to be a differentiable manifold (or a differentiable manifold with boundary), although they are topological manifolds with boundaries. For $k>3$, the geometry is more complicated, as $\mathcal{M}_{k}^{\varepsilon}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ decompose into strata of different codimensions. In particular, understanding the structure of the collision manifold is of independent interest from a geometric perspective.
Therefore, in order to carry out Riemannian geometric considerations such as defining tangent and normal vectors, we restrict our analysis to the smooth non-collision manifold $\mathcal{M}_{k}^{\varepsilon,\geq d}$ by removing the collision region $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$.
At each point $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ on the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq d}_{k}$, we define the normal space to the manifold as follows
$V_{\xi_{1},\dots,\xi_{k}}=\big\{v\in L^{2}:\;\langle v,\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle=0\quad\mbox{for all }1\leq j\leq k\big\}.$ (4.11)
Here the normal space $V_{\xi_{1},\dots,\xi_{k}}$ at $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ is a subspace of codimension $k$ in $L^{2}$, orthogonal to the tangent vectors
$\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\dots,\partial_{\xi_{k}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$
of the multi-soliton manifold $\mathcal{M}^{\varepsilon}_{k}$. Thanks to Lemma 3.1-(iii), each tangent vector $\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ is highly localized around its center $\xi_{j}$ with an exponentially decaying tail. Therefore,
$|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\partial_{\xi_{i}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle|\lesssim e^{-c|\xi_{i}-\xi_{j}|}$ (4.12)
for $i\neq j$. This implies that the tangent vectors are almost orthogonal when they do not collide, that is, when $\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$.
We are now ready to define the projection map $\pi^{\varepsilon}$ onto the multi-soliton manifold, introduced in Theorem 1.4. Let $\mathcal{M}$ be a compact manifold in a Hilbert space $\mathcal{H}$. If $\delta>0$ is sufficiently small, we can assign to any $\phi\in\mathcal{H}$ with $\text{dist}(\phi,\mathcal{M})<\delta$ a unique closest point $\pi(\phi)$ in the manifold $\mathcal{M}$. This follows from the $\varepsilon$-neighborhood theorem [24, p.69]. If $\text{dist}(\phi,\mathcal{M})\geq\delta$, then we set $\pi(\phi)=0$. Recall that
$\mathcal{M}^{\varepsilon,\geq d}_{k}:$ $=\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\}$
is a $k$-dimensional compact manifold. Therefore, if a field $\phi$ satisfies
$\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})=\inf_{\begin{subarray}{c}-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\end{subarray}}\|\phi-m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}<\delta$ (4.13)
for sufficiently small $\delta>0$, we can assign a unique pair $(\xi_{1},\dots,\xi_{k})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ such that
$\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}+v,$ (4.14)
where the normal coordinate $v\in V_{\xi_{1},\dots,\xi_{k}}$ satisfies $\|v\|_{L^{2}}<\delta$. Therefore, according to (4.14), we are now able to define the projection $\pi^{\varepsilon}$ onto the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq d}_{k}$ as the closest point
$\pi^{\varepsilon}(\phi)=\begin{cases}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})<\delta\\ 0,\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})\geq\delta.\end{cases}$ (4.15)
Remark 4.2.
For sufficiently large $d>0$, we can define a projection map onto the (non-approximating) manifold
$\mathcal{M}^{\geq d}_{k}:$ $=\{m_{\xi_{1},\dots,\xi_{k}}:-\infty<\xi_{1}\leq\dots\xi_{k}<\infty\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\}$
with a uniform neighborhood of size $\delta>0$ since the Jacobian matrix has the uniform lower bound
$\textup{det}(\langle\partial_{\xi_{i}}m_{\xi_{1},\dots,\xi_{k}},\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle_{1\leq i,j\leq k})\gtrsim\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{k}(1+O(e^{-cd}))$
as $d\to\infty$. Note that this allows us to apply the implicit function theorem with uniform control.
We now introduce a disintegration formula from [19, Lemma 3], which expresses Gaussian functional integrals on a small neighborhood of the compact manifold $\mathcal{M}^{\varepsilon,\geq d}_{k}$, defined in (4.10), in terms of tangential $\xi_{1},\dots,\xi_{k}$ and normal coordinates $v$.
Lemma 4.3.
Let $F$ be a bounded, continuous function on $L^{2}$. Then, we have
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{k}^{\varepsilon,\geq d})<\delta\}}F(\phi)\mu_{\varepsilon}^{k}(d\phi)$
$=\int\dots\int\limits_{U_{\varepsilon}}F(m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}+v)e^{-\frac{1}{2}\|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}^{2}-\langle(-\partial_{x}^{2})m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},v\rangle_{L^{2}}}$
$\hphantom{XXXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v)\,\mu^{\perp}_{\varepsilon,\xi_{1},\dots,\xi_{k}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{k}),$ (4.16)
where $\mathcal{M}^{\varepsilon,\geq d}_{k}$ is as defined in (4.10), $\Delta_{k}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{k}\leq\overline{L}_{\varepsilon}\}$,
$U_{\varepsilon}=\big\{(\xi_{1},\dots,\xi_{k},v)\in\Delta_{k}\times V_{\xi_{1},\dots,\xi_{k}}:\|v\|_{L^{2}}<\delta\;\textup{and}\;\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\big\}$
$\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v):=\det\big(\operatorname{Id}-W_{\xi_{1},\dots\xi_{k},v}\big),$ (4.17)
and $\mu^{\perp}_{\varepsilon,\xi_{1},\dots,\xi_{k}}$ is the Gaussian measure on the normal space $V_{\xi_{1},\dots,\xi_{k}}$ with covariance $\varepsilon(-\partial_{x}^{2})^{-1}$, subject to Dirichlet boundary conditions on $L_{\varepsilon},L_{\varepsilon}]$
$\mu^{\perp}_{\varepsilon,\xi_{1},\dots,\xi_{k}}(dv)=Z_{\varepsilon,\xi_{1},\dots,\xi_{k}}^{-1}e^{-\frac{1}{2\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}v|^{2}dx}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x).$
In (4.17) $W_{\xi_{1},\dots,\xi_{k},v}$ denotes the Weingarten map, defined in (4.20). In addition, $d\sigma$ is the surface measure on the manifold, parametrized by $(\xi_{1},\dots,\xi_{k})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, as defined in (4.19).
Proof.
The formula (4.16) follows from [19, Lemma 3]. ∎
Remark 4.4.
The disintegration formula has been widely employed in recent works [33, 36, 22] in settings where the energy functional admits explicit minimizers and the analysis is carried out around a single-soliton manifold. However, to the best of our knowledge, in Proposition 7.4 this is the first time the disintegration formula is applied in a regime where no minimizers exist and the analysis is performed around a multi-soliton manifold.
Note that the geometry of the smooth soliton manifold $\mathcal{M}^{\varepsilon,\geq d}_{k}$ is reflected in the surface measure $d\sigma$ and the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}$. The orthonormal vectors $t_{j}=t_{j}(\xi_{1},\dots,\xi_{k})$, for $j=1,\dots,k$, are obtained by applying the Gram-Schmidt orthonormalization procedure in $L^{2}$ to the tangent vectors
$\big\{\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\dots,\partial_{\xi_{k}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\big\}$
as follows
$t_{j}:=\frac{w_{j}}{\|w\|_{L^{2}}}\qquad\textup{and}\qquad w_{j}:=\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}-\sum_{\ell<j}\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle_{L^{2}}\;t_{\ell}.$ (4.18)
Then, the surface measure $d\sigma(\xi_{1},\dots,\xi_{k})$ is given by
$d\sigma(\xi_{1},\dots,\xi_{k})=|\gamma(\xi_{1},\dots,\xi_{k})|d\xi_{1}\dots d\xi_{k},$ (4.19)
where
$\gamma(\xi_{1},\dots,\xi_{k})=\det\begin{pmatrix}\langle\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{1}\rangle_{L^{2}}&\cdots&\langle\partial_{\xi_{k}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{1}\rangle_{L^{2}}\\ \vdots&\ddots&\vdots\\ \langle\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{k}\rangle_{L^{2}}&\cdots&\langle\partial_{\xi_{k}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{k}\rangle_{L^{2}}\end{pmatrix}$.
Indeed, thanks to (4.12), the tangent vectors $\{\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\}_{j=1}^{k}$ are almost orthogonal, with an error of order $e^{-d}$, where $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d$. Hence, the Jacobian matrix is nearly diagonal, with diagonal entries given by $\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}$,
$|\gamma(\xi_{1},\dots,\xi_{k})|=k\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd})),$
uniformly in $\xi_{1},\dots\xi_{k}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\to\infty$.
In (4.17) the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}$ encodes the curvature of the surface $\mathcal{M}^{\varepsilon,\geq d}_{k}$ by capturing how the normal vector $v$ changes direction as we move along different tangent directions on the surface. More precisely, the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}=-dN_{\xi_{1},\dots,\xi_{k}}(v)$ at a point $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\in\mathcal{M}^{k,\varepsilon}$, defined via the differential of the Gauss map $N_{\xi_{1},\dots,\xi_{k}}$ at $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\in\mathcal{M}^{\varepsilon,\geq d}_{k}$, is the linear map
$W_{\xi_{1},\dots,\xi_{k},v}:T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}\to T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k},$ (4.20)
where $T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}:=\text{span}\big\{t_{1},\dots,t_{k}\big\}$ is the tangent space of $\mathcal{M}^{\varepsilon,\geq d}_{k}$ at $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ and $t_{1},\dots,t_{k}$ are obtained from the Gram–Schmidt orthonormalization described above (see (4.18)). Specifically, the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}$ in the basis $\{t_{1},\dots,t_{k}\}$ is given by
$\begin{pmatrix}\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{k}\rangle_{L^{2}}\\ \vdots&\ddots&\vdots\\ \langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{k}\rangle_{L^{2}}\end{pmatrix}$,
where
$L^{2}\ni v\mapsto N_{\xi_{1},\dots,\xi_{k}}(v)$
is a parametrization of the normal space to $\mathcal{M}^{\varepsilon,\geq d}_{k}$ at a point $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$. In particular, the $k\times k$ determinant
$\det\big(\operatorname{Id}-W_{\xi_{1},\dots,\xi_{k},v}\big)=1+O\big(\|v\|_{L^{\infty}}^{k}\big)$ (4.21)
is a $k$-th order function of $v$.
Before concluding this subsection, we present the following lemma, which will be used later.
Lemma 4.5.
Let $t_{1},\dots,t_{k}$ be obtained by applying the Gram–Schmidt orthonormalization in $L^{2}$ to $\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\dots,\partial_{\xi_{k}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$, as described in (4.18). Then, for each $j$, we have
$|t_{j}(x)|\lesssim e^{-c|x-\xi_{j}|}$
provided that $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d$ for $d$ sufficiently large.
Proof.
Note that $t_{1}$ is localized at $\xi_{1}$ with an exponentially decaying tail. Inductively assume $t_{\ell}$ are localized at $\xi_{\ell}$ for $\ell<j$. Then $|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle|\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}$ and so
$|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle t_{\ell}(x)|\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}e^{-c_{1}|x-\xi_{\ell}|}\lesssim e^{-c\min_{j\neq\ell}|\xi_{j}-\xi_{\ell}|}.$
This, together with (4.18), implies that
$|w_{j}(x)|$ $\leq|\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}(x)|+\sum_{\ell<j}|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle_{L^{2}}||t_{\ell}(x)|\lesssim e^{-c|x-\xi_{j}|}$
as $\min_{j\neq\ell}|\xi_{j}-\xi_{\ell}|\geq d\to\infty$. ∎
5. Gaussian measures associated with Schrödinger operators
In this section, we study the linearized operator around the multi-soliton configuration. For well-separated centers $\xi_{1},\dots,\xi_{k}$ satisfying
$\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\to\infty$
as $\varepsilon\to 0$, the Hessian of the energy at $m_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m_{\xi_{j}}$ is
$\nabla^{2}E(m_{\xi_{1},\dots,\xi_{k}})$ $=-\partial_{x}^{2}+\cos(m_{\xi_{1},\dots,\xi_{k}})$
$=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\textup{sech}^{2}(\cdot-\xi_{j})+O(e^{-c\min_{i\neq j}|\xi_{i}-\xi_{j}|}),$ (5.1)
where the second equality follows from the well-separated condition $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\to\infty$ and the structure of the multi-soliton. Note that the error term is exponentially small in the minimal separation $d_{\varepsilon}$ and uniform in all $\xi_{1},\dots,\xi_{k}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$. In Section 7, $\nabla^{2}E(m_{\xi_{1},\dots,\xi_{k}})$ plays the role of the covariance operator for a new Gaussian measure.
We begin by recalling the following well-known spectral properties of the linearized operator $\nabla^{2}E(m_{\xi})=\mathcal{L}_{\xi}$ around the single kink
$\mathcal{L}_{\xi}=-\partial_{x}^{2}+\cos(m_{\xi})=-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi),$ (5.2)
where $\xi\in\mathbb{R}$. The potential is of the reflectionless Pöschl–Teller type.
Lemma 5.1.
Let $\xi\in\mathbb{R}$.
  • (1)
    The linearized operator $\mathcal{L}_{\xi}$, defined in (5.2), is self-adjoint. Its spectrum is given by
    $\sigma(\mathcal{L}_{\xi})=\sigma_{d}\cup\sigma_{c}=\{0\}\cup[1,\infty).$
  • (2)
    The eigenfunction corresponding to the zero eigenvalue arises from the tangent vector $\partial_{\xi}m_{\xi}$ to the soliton manifold $\{m(\cdot-\xi)\}_{\xi\in\mathbb{R}}$, which is associated with translation invariance
    $\mathcal{L}_{\xi}(\partial_{\xi}m_{\xi})=0.$ (5.3)
  • (3)
    The linearized operator $\mathcal{L}_{\xi}$ satisfies the following coercivity: there exists $\lambda_{0}\in[1,\infty)$ such that
    $\langle\mathcal{L}_{\xi}v,v\rangle\geq\lambda_{0}\|v\|_{H^{1}}^{2}.$ (5.4)
    for every $v\in H^{1}(\mathbb{R})$ with $\langle v,\partial_{\xi}m_{\xi}\rangle_{L^{2}(\mathbb{R})}=0$.
Proof.
For the proof, see [15, Lemma 3.1]. ∎
We denote
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}:=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j}).$
Note that the tangent vectors with respect to the center coordinates are
$\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}=\partial_{\xi_{j}}m_{\xi_{j}}=-2\text{sech}(\cdot-\xi_{j}),\quad j=1,\dots,k.$
This implies that
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}})$ $=(-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j}))\partial_{\xi_{j}}m_{\xi_{j}}+4\sum_{i\neq j}\text{sech}^{2}(\cdot-\xi_{i})\text{sech}(\cdot-\xi_{j})$
$=4\sum_{i\neq j}\text{sech}^{2}(\cdot-\xi_{i})\text{sech}(\cdot-\xi_{j}),$
where we used $\mathcal{L}_{\xi_{j}}(\partial_{\xi_{j}}m_{\xi_{j}})=0$ from (5.3). Hence, the linearized operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ satisfies
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}),\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle$ $=-8\sum_{i\neq j}\int_{\mathbb{R}}\text{sech}^{2}(x-\xi_{i})\text{sech}^{2}(x-\xi_{j})dx$
$=-O(e^{-\min_{i\neq j}|\xi_{i}-\xi_{j}|}),$
In the second line, we used the fact that $\text{sech}(\cdot-\xi_{i})$ is localized around $\xi_{i}$ with exponentially decaying tails. Consequently, $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ has $k$ near-zero (negative) eigenvalues of size $O(e^{-cd_{\varepsilon}})$, where $d_{\varepsilon}=\min_{i\neq j}|\xi_{i}-\xi_{j}|$.
The next lemma shows that if the centers of the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}$ are sufficiently separated, then the linearized operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ is uniformly coercive on the normal space, that is, after projecting off the tangent vectors.
Lemma 5.2.
There exists $\lambda_{0},d_{0}>0$ such that, under the separation condition $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{0}$, the coercivity
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}v,v\rangle_{L^{2}(\mathbb{R})}\geq\lambda_{0}\|v\|_{H^{1}(\mathbb{R})}^{2}$
holds for every $v\in H^{1}(\mathbb{R})$ with $\langle v,\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle_{L^{2}(\mathbb{R})}=0$, $j=1,\dots,k$.
Proof.
Choose a smooth partition of unity such that $\sum_{j=0}^{k+1}\chi_{j}^{2}=1$ and $\sup_{j}\|\partial_{x}\chi_{j}\|_{L^{\infty}}\lesssim d_{\varepsilon}^{-1}$, where $d_{\varepsilon}=\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$ as $\varepsilon\to 0$. For $j=1,\dots,k$, $\chi_{j}$ is supported in a fixed neighborhood of $\xi_{j}$, and $\chi_{0}$, $\chi_{k+1}$ cover the left/right far field (vacua 0 and $2\pi Q$). Then, we have
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}v,v\rangle_{L^{2}(\mathbb{R})}$ $=\sum_{j}\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-\sum_{j}\|(\partial_{x}\chi_{j})v\|_{L^{2}}^{2}$
$\geq\sum_{j}\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Cd_{\varepsilon}^{2}\|v\|_{L^{2}}^{2},$ (5.5)
where the term $\sum_{j=1}^{k}\|(\partial_{x}\chi_{j})v\|_{L^{2}}^{2}$ comes from expanding the kinetic term after applying a partition of unity. We analyze each localized piece to obtain a uniform coercivity estimate. Note that
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v)$ $=\mathcal{L}_{\xi_{j}}(\chi_{j}v)-2\sum_{i:i\neq j}\text{sech}^{2}(\cdot-\xi_{i})\chi_{j}v$
$=\mathcal{L}_{\xi_{j}}(\chi_{j}v)+O(e^{-cd_{\varepsilon}})\cdot\chi_{j}v,$
where $\mathcal{L}_{\xi_{j}}=-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j})$. This implies that
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Ce^{-cd_{\varepsilon}}\|\chi_{j}v\|_{L^{2}}^{2}.$ (5.6)
We now show that $\chi_{j}v$ is almost orthogonal to $\partial_{\xi_{j}}m_{\xi_{j}}$ if $\langle v,\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle=0$. Indeed,
$\langle\chi_{j}v,\partial_{\xi_{j}}m_{\xi_{j}}\rangle_{L^{2}(\mathbb{R})}$ $=\langle v,\partial_{\xi_{j}}m_{\xi_{j}}\rangle_{L^{2}(\mathbb{R})}-\sum_{i:i\neq j}\langle v,\chi_{i}(\partial_{\xi_{j}}m_{\xi_{j}})\rangle_{L^{2}(\mathbb{R})}$
$=-\sum_{i:i\neq j}\langle v,\chi_{i}(\partial_{\xi_{j}}m_{\xi_{j}})\rangle_{L^{2}(\mathbb{R})},$
where we used $\langle v,\partial_{\xi_{j}}m\rangle=\langle v,\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle=0$. Since $\partial_{\xi_{j}}m_{\xi_{j}}$ is localized around $\xi_{j}$ and has exponentially decaying tail, we obtain almost orthogonality
$|\langle\chi_{j}v,\partial_{\xi_{j}}m_{\xi_{j}}\rangle_{L^{2}(\mathbb{R})}|\lesssim\sum_{i:i\neq j}e^{-cd_{\varepsilon}}\|\chi_{i}v\|_{L^{2}}^{2}.$ (5.7)
Combining the almost-orthogonality (5.7) with the coercivity estimate (5.4) for the linearized operator $\mathcal{L}_{\xi_{j}}$ around the single kink $m_{\xi_{j}}$, we obtain
$\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\frac{\lambda_{0}}{2}\|\chi_{j}v\|_{H^{1}}^{2}-Ce^{-cd_{\varepsilon}}\sum_{i:i\neq j}\|\chi_{i}v\|_{L^{2}}^{2}$ (5.8)
for some $\lambda_{0}>0$. Putting the pieces together with (5.5), (5.6), (5.8), and taking $d_{\varepsilon}\geq d_{0}$ large enough, we obtain
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}v,v\rangle_{L^{2}(\mathbb{R})}$ $\geq\sum_{j}\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Cd_{\varepsilon}^{2}\|v\|_{L^{2}}^{2}$
$\geq\frac{\lambda_{0}}{2}\sum_{j=1}^{k}\|\chi_{j}v\|_{H^{1}}^{2}-C(e^{-cd_{\varepsilon}}+d_{\varepsilon}^{2})\|v\|_{L^{2}}^{2}$
$\geq\frac{\lambda_{0}}{4}\|v\|_{H^{1}}^{2}.$
This completes the proof of Lemma 5.2.
From now on, we study the operator
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
on the finite interval $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions.
Lemma 5.3.
Let $d_{0}>0$ be as in Lemma 5.2. Then there exits $L_{0},\zeta>0$ such that if $L_{\varepsilon}\geq L_{0}$ and $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{0}$,
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}v,v\rangle_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}\geq\zeta\|v\|^{2}_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}$
holds for every $v\in H^{1}_{0}([-L_{\varepsilon},L_{\varepsilon}])$ with $\langle v,\partial_{\xi_{j}}m_{\xi_{1},\dots,\xi_{k}}\rangle_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}=0$, $j=1,\dots,k$. Here $\zeta$ is independent of $L_{\varepsilon},\xi_{1},\dots,\xi_{k}$.
Proof.
Assume not. Then there exist sequences $L_{n}\to\infty$, configurations $\xi_{1}^{n},\dots,\xi_{k}^{n}$ with $\min_{i\neq j}|\xi_{i}^{n}-\xi_{j}^{n}|\geq d_{0}$, and $v_{n}\in H^{1}_{0}([-L_{n},L_{n}])$ with $\langle v_{n},\partial_{\xi_{j}^{n}}m_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\rangle_{L^{2}([-L_{n},L_{n}])}=0$ such that
$\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}v_{n},v_{n}\rangle_{L^{2}([-L_{n},L_{n}])}\leq\frac{1}{n}\|v_{n}\|^{2}_{L^{2}([-L_{n},L_{n}])}.$
We extend $v_{n}$ by zero outside $L_{\varepsilon},L_{\varepsilon}]$ and denote the resulting function by $\widetilde{v}_{n}$. Then $\widetilde{v}_{n}\in H^{1}(\mathbb{R})$ and
$\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}=\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}v_{n},v_{n}\rangle_{L^{2}([-L_{n},L_{n}])}\leq\frac{1}{n}\|v_{n}\|^{2}_{L^{2}([-L_{n},L_{n}])}=\frac{1}{n}\|\widetilde{v}_{n}\|^{2}_{L^{2}(\mathbb{R})},$ (5.9)
where we used the Dirichlet boundary condition. Since
$\langle\widetilde{v}_{n},\partial_{\xi_{j}^{n}}m_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\rangle_{L^{2}(\mathbb{R})}=\langle v_{n},\partial_{\xi_{j}^{n}}m_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\rangle_{L^{2}([-L_{n},L_{n}])}=0$
for $j=1,\dots,k$, Lemma 5.2 implies that
$\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}\geq\lambda_{0}\|\widetilde{v}_{n}\|_{L^{2}(\mathbb{R})}^{2}.$ (5.10)
Combining (5.9) and (5.10) gives a contradiction for large $n$. Hence, a uniform $\zeta>0$ exists.
According to Lemma 5.3, to avoid zero eigenvalues, we need to project off the tangential directions. Once the zero modes are removed, we can invert the operator on the subspace, namely, the normal space $V_{\xi_{1},\dots,\xi_{k}}$. We define the projected operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$ (5.11)
viewed as an operator on the finite interval $L_{\varepsilon},L_{\varepsilon}]$, with Dirichlet boundary conditions. Here the projection $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ is given by
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\langle\cdot,t_{j}\rangle t_{j}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j}$ (5.12)
where $t_{1},\dots,t_{k}$ are obtained by applying the Gram–Schmidt orthonormalization to the tangent vectors $\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$, $j=1,\dots,k$, as described above (4.18), and
$\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}.$
As mentioned in Section 2, the inner product $\langle\cdot,\cdot\rangle$ is understood as $\langle\cdot,\cdot\rangle_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}$ unless otherwise specified.
We are now ready to define the Gaussian measure with the covariance operator $C_{\xi_{1},\dots,\xi_{k}}$.
Lemma 5.4.
There exist sufficiently large $L_{0}$, $d_{0}>0$ such that if $L_{\varepsilon}\geq L_{0}$ and $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{0}$, we can define the Gaussian measure
$d\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}=Z_{\xi_{1},\dots,\xi_{k}}^{-1}e^{-\frac{1}{2}\langle C_{\xi_{1},\dots,\xi_{k}}^{-1}v,v\rangle}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x).$
for any fixed $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]$.
Proof.
The operator
$-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
considered on the finite interval $L_{\varepsilon},L_{\varepsilon}]$, is a self-adjoint Sturm–Liouville operator since $2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$ is a Schwartz function. Hence, the operator has purely discrete spectrum $\{\lambda_{n}\}_{n}$ with $\lambda_{n}\sim|n|^{2}$ as $n\to\infty$. In particular, if the operator is strictly positive, its inverse belongs to the trace class. By Lemma 5.3, the projected operator is strictly positive, hence its inverse on that subspace is trace class. Therefore, the corresponding Gaussian measure is well defined.
Before concluding this subsection, we present the structure of the partition function of the Gaussian measure in Lemma 5.4. As the interaction between solitons becomes negligible, the spectrum of the multi-well operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ becomes exactly $k$ copies of the single-well spectrum.
Lemma 5.5.
Let $Q\in\mathbb{Z}$ with $|Q|=k$ and $d_{\varepsilon}=|\log(\varepsilon\log\frac{1}{\varepsilon})|$. Then we have
$Z_{\xi_{1},\dots,\xi_{k}}=(Z_{0})^{k}(1+O(e^{-cd_{\varepsilon}})),$
uniformly in $\xi_{1},\dots,\xi_{k}$ satifying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$, where $Z_{0}$ is the partition function of the Gaussian measure associated with the single-well operator $\mathcal{L}_{0}=-\partial_{x}^{2}+1-\textup{sech}^{2}(x)$.
Proof.
Since $\mathcal{L}_{0}=-\partial_{x}^{2}+1-\textup{sech}^{2}(x)$ acts on the normal space and is a Sturm–Liouville operator on the compact interval $L_{\varepsilon},L_{\varepsilon}]$, it has a discrete spectrum
$1\leq\lambda_{1}<\lambda_{2}<\dots<\lambda_{\ell}<\dots\to\infty.$ (5.13)
Let $\{\phi\}_{\ell\geq 1}$ be an orthonormal eigenbasis of $\mathcal{L}_{0}$ on the normal space with eigenvalues $\lambda_{\ell}\geq 1$, thanks to Lemma 5.1. Since $\mathcal{L}_{0}$ is a bounded perturbation of $\partial_{x}^{2}+1$ by $\textup{sech}^{2}$, we can show that
$|\phi_{\ell}(x)|+|\partial_{x}\phi_{\ell}(x)|\lesssim e^{-c|x|}.$ (5.14)
For each $\xi_{j}$, $j=1,\dots,k$, we define
$\phi_{\ell,j}(x):=\phi_{\ell}(x-\xi_{j}).$
We can easily verify that $\phi_{\ell,j}$ is an eigenfunction of $\mathcal{L}_{\xi_{j}}=-\partial_{x}^{2}+1-\textup{sech}^{2}(\cdot-\xi_{j})$ with the same eigenvalue $\lambda_{\ell}$. Furthermore the translated eigenfunctions $\phi_{\ell,j}$ act as approximate eigenfunctions for $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\textup{sech}^{2}(\cdot-\xi_{j})$ as follows
$\langle\phi_{\ell,j},\mathcal{L}_{\xi_{1},\dots,\xi_{k}}\phi_{\ell,j}\rangle$ $=\langle\phi_{\ell,j},(-\partial_{x}^{2}+1-2\textup{sech}^{2}(\cdot-\xi_{j}))\phi_{\ell,j}\rangle-2\sum_{i\neq j}\int\textup{sech}^{2}(x-\xi_{i})|\phi_{\ell,j}(x)|^{2}dx$
$=(\lambda_{\ell}+O(e^{-cd_{\varepsilon}}))\|\phi_{\ell,j}\|_{L^{2}}^{2},$
where we used (5.14). Hence, every single-well level $\lambda_{\ell}$ generates a cluster of $k$ nearby eigenvalues $\lambda_{\ell,1},\dots,\lambda_{\ell,k}$ of the multiwell operator $\mathcal{L}_{\xi_{1}\dots,\xi_{k}}$
$\lambda_{\ell,j}=\lambda_{\ell,j}(\xi_{1},\dots,\xi_{k})=\lambda_{\ell}+O(e^{-cd_{\varepsilon}}),\quad j=1,\dots,k.$ (5.15)
By using the fact that $\{\phi_{\ell,j}\}_{j=1,\dots,k}$ is an almost-orthonormal family when $d_{\varepsilon}$ is sufficiently large, and there exists a single-well spectral gap $\lambda_{{}_{\ell}+1}-\lambda_{\ell}>0$ in (5.13), we can easily check that each cluster has exactly $k$ members and the clusters are disjoint.
To describe all discrete eigenvalues for the multi-well operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$, we take the product over all $\ell\geq 1$ and use (5.15). More precisely, for any finite $N$,
$\textup{det}_{V_{\xi_{1},\dots,\xi_{k}}}(\mathbf{P}_{N}\mathcal{L}_{\xi_{1},\dots,\xi_{k}}\mathbf{P}_{N})$ $=\prod_{\ell=1}^{N}\bigg(\prod_{j=1}^{k}\lambda_{\ell,j}\bigg)=\bigg(\prod_{\ell=1}^{N}\lambda_{\ell}\bigg)^{k}(1+O(e^{-cd_{\varepsilon}}))$
$=\textup{det}_{V_{0}}(\mathbf{P}_{N}\mathcal{L}_{0}\mathbf{P}_{N})^{k}(1+O(e^{-cd_{\varepsilon}})),$
uniformly in $\xi_{1},\dots,\xi_{k}$ satifying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$, where $\mathbf{P}_{N}$ is the finite-dimensional projection onto modes with $|\ell|\leq N$.
6. Concentration around the multi-soliton manifold without collisions
In this section, we prove Theorem 1.2.
6.1. Free energy
Before proving Theorem 1.2, we study in this subsection the behavior of the free energy as $\varepsilon\to 0$.
Proposition 6.1.
Let $Q\in\mathbb{Z}$ and $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$. Then, we have
$\lim_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}=-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi),$
where $\mathcal{C}_{Q}$ is the topological sector in (1.3) and $Z^{Q}_{\varepsilon}$ is the partition function of the Gibbs measure $\rho^{Q}_{\varepsilon}$ defined in (1.5).
Remark 6.2.
Based on Lemma 2.2 (especially (2.8)), a naive calculation would restrict the interval size to $L_{\varepsilon}=\varepsilon^{-\frac{1}{3}+\eta}$. However, by exploiting the specific structure of the sine–Gordon model, we are able to extend the interval up to $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$. See (6.6), (6.7), and (6.8).
Proof.
By using the Boué–Dupuis formula, Lemma 2.1, we write
$\log Z^{Q}_{\varepsilon}$ $=\log\mathbb{E}_{\mu_{\varepsilon}^{Q}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\bigg]$
$=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\big(1-\cos(\ell^{Q}+\sqrt{\varepsilon}Y+\sqrt{\varepsilon}\Theta)\big)dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg],$
where we used the notation $Y:=Y(1)$ and $\Theta:=\Theta(1)$ for convenience.
We first derive an upper bound on the free energy. By performing the change of variables $\sqrt{\varepsilon}\theta\to\theta$, which does not affect the variational problem, we can rewrite
$\varepsilon\log Z^{Q}_{\varepsilon}$ $=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\sqrt{\varepsilon}\Theta))dx-\frac{\varepsilon}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg]$
$=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg]$
$\leq\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg],$ (6.1)
where in the last line we used (2.6) and $\dot{\mathbb{H}}^{1}$ denotes the space of $\dot{H}^{1}$-valued random variables. We now take the change of variable
$W=\ell^{Q}+\Theta.$ (6.2)
Combining (6.1) and (6.2) yields
$\varepsilon\log Z_{\varepsilon}^{Q}$ $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1}{2}\|\partial_{x}W\|_{L^{2}}^{2}$
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg].$ (6.3)
By using Young’s inequality, for any small $\delta>0$ we have
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg|\leq\frac{\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2},$ (6.4)
where $C_{\delta}$ is a large constant. By plugging (6.4) into (6.3), we have
$\varepsilon\log Z_{\varepsilon}^{Q}$ $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1-\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}\bigg]+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}.$ (6.5)
By taking the Taylor expansion around $W$, we have
$\cos(\sqrt{\varepsilon}Y+W)=\cos(W)-\sin(W)\cdot(\sqrt{\varepsilon}Y)+O\big((\sqrt{\varepsilon}Y)^{2}\big).$ (6.6)
Note that $\sin^{2}(W)=(1-\cos(W))(1+\cos(W))\leq 2(1-\cos(W))$. Hence,
$|\sin(W)\cdot(\sqrt{\varepsilon}Y)|\leq\kappa(1-\cos(W))+C_{\kappa}(\sqrt{\varepsilon}Y)^{2}$ (6.7)
for an arbitrarily small $\kappa>0$ and a large constant $C_{\kappa}$. By plugging (6.6) into (6.5) and using (6.7), we obtain
$\varepsilon\log Z^{Q}_{\varepsilon}$ $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg]$
$\hphantom{X}+\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}C_{\kappa}|\sqrt{\varepsilon}Y|^{2}dx\bigg]+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}.$
Using Lemma 2.2 and $\partial_{x}\ell^{Q}=\frac{\pi Q}{L_{\varepsilon}}$,
$\varepsilon\log Z_{\varepsilon}^{Q}$
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg]+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}.$ (6.8)
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\delta\to 0$, and $\kappa\to 0$ in order, we obtain
$\limsup_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\leq-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
From now on, we study the lower bound on the free energy. From (6.1), recall that
$\varepsilon\log Z_{\varepsilon}^{Q}=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg].$ (6.9)
We choose a drift $\theta=\theta^{0}$ by
$\theta^{0}(t)=\eta^{-1}\cdot\mathbf{1}_{\{t>1-\eta\}}(t)(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),$ (6.10)
where $\eta>0$, $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]$, and
$Y_{N}(1-\eta)=\sum_{1\leq n\leq N}\frac{B_{n}(1-\eta)}{\sqrt{\lambda_{n}}}e_{n}(x),$
approximating the Gaussian field $Y=Y(1)=\sum_{n\geq 1}\frac{B_{n}(1)}{\sqrt{\lambda_{n}}}e_{n}(x)$. Then the definition of $\Theta$ in (2.6) implies that
$\Theta^{0}:=\int_{0}^{1}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta^{0}(t)dt=-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}.$ (6.11)
In (6.10) the choice of the drift $\theta^{0}(t)$ is admissible $\theta^{0}\in\mathbb{H}_{a}$ since it is an adapted process due to the cutoff $\mathbf{1}_{\{t>1-\eta\}}$ and it satisfies the required regularity condition thanks to the truncation to the modes $1\leq n\leq N$. By plugging (6.10) and (6.11) into (6.9), we obtain
$\varepsilon\log Z_{\varepsilon}^{Q}\geq\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))dx-\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\bigg].$ (6.12)
From the definition of $\theta^{0}$ in (6.10), we write
$\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt=\frac{1}{2}\|(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\|_{L^{2}_{x}}^{2}.$
Using the inequality $(a+b+c)^{2}\leq(1+\delta)|a|^{2}+C_{\delta}(|b|^{2}+|c|^{2})$ for any real numbers $a,b,c$, and $\delta>0$, which follows from Young’s inequality, we have
$\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\leq\frac{1+\delta}{2}\|\partial_{x}m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon}\|_{L^{2}}^{2}+\frac{C_{\delta}}{2}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\eta))\|_{L^{2}}^{2}+\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}\big)$ (6.13)
for any $\delta>0$ and the corresponding constant $C_{\delta}>0$. Regarding the potential energy term, by taking the Taylor expansion around $m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$,
$\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))$
$=\cos(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})-\sin(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})\cdot\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+O\big((\sqrt{\varepsilon}(Y-Y_{N}(1-\eta)))^{2}\big).$ (6.14)
By plugging (6.14) and (6.13) into (6.12), we obtain
$\varepsilon\log Z^{Q}_{\varepsilon}$ $\geq\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})dx-\frac{1+\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon}|^{2}dx\bigg]$
$\hphantom{X}-\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))|^{2}dx\bigg]$
$\hphantom{X}-C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}-C_{\delta}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\eta))\|_{L^{2}}^{2},$ (6.15)
where we used $\mathbb{E}\big[\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))\big]=0$. Using Lemma 2.2 and $\partial_{x}\ell^{Q}=\frac{\pi Q}{L_{\varepsilon}}$, we obtain
$\varepsilon\log Z^{Q}_{\varepsilon}$ $\geq\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})dx-\frac{1+\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}|^{2}dx\bigg]$
$\hphantom{X}-c\varepsilon L_{\varepsilon}^{2}-C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}-C_{\delta}\varepsilon\|\partial_{x}Y_{N}(1-\eta)\|_{L^{2}}^{2}.$
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$ and $\delta\to 0$ in order, we obtain
$\liminf_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\geq-E(m_{\xi_{1},\dots,\xi_{Q}}).$
By letting $\min_{i\neq j}|\xi_{i}-\xi_{j}|\to\infty$, and using Remark 3.5, we obtain
$\liminf_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\geq-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
This completes the proof of Proposition 6.1.
In order to prove the main result of this section, that is, the large deviation estimates, we need the following lemma.
Lemma 6.3.
Let $Q\in\mathbb{Z}$ with $Q>0$ and $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$. Then the following holds.
  • (i)
    Let $\delta>0$. Then we have
    $\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi).$
  • (ii)
    Let $d,\delta>0$. Then
    $\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\end{subarray}}E(\phi).$
Proof.
As in (6.1), using the Boué–Dupuis formula (Lemma 2.1) and the change of variables $\sqrt{\varepsilon}\theta\to\theta$, we write
$\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta\}}\bigg]$
$\leq\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\cdot\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta\}}\Big\}\bigg]$
$\leq\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx\cdot\mathbf{1}_{\{\textup{dist}(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta,\mathcal{M}^{Q})\geq\delta\}}-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg].$ (6.16)
By taking the change of variable $W=\ell^{Q}+\Theta$, we write
[…]
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg]$
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[\bigg(-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1}{2}\|\partial_{x}W\|_{L^{2}}^{2}\bigg)\cdot\mathbf{1}_{\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q})\geq\delta\}}$
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg].$ (6.17)
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, we have $\mathbb{E}\big[\|\sqrt{\varepsilon}Y\|_{L^{2}}^{2}\big]\to 0$ by (2.8). Therefore, with high probability
$\{\textup{dist}(W,\mathcal{M}_{Q})\geq 2\delta\}\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q})\geq\delta\}.$ (6.18)
By following the steps used to obtain (6.8) together with (6.18), we obtain
[…]
$\hphantom{XXXXXXXXX}+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\eta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}$ (6.19)
for any $\eta>0$ and $\kappa>0$. This implies that under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\eta\to 0$, and $\kappa\to 0$ in order, we obtain
$\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi).$
This proves the first part of Lemma 6.3.
For the second part of Lemma 6.3, we only need to modify the following part. Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, we have $\mathbb{E}\big[\|\sqrt{\varepsilon}Y\|_{L^{2}}^{2}\big]\to 0$ by (2.8). Therefore, with high probability
$\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\}$ $\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q}^{<d})<\delta\}.$ (6.20)
By proceeding as in obtaining (6.19), together with (6.20), we obtain
$\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{<d}_{Q})<\delta\}}\bigg]$
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[\bigg(-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\eta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg)\mathbf{1}_{\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\}}\bigg]$
$\hphantom{XXXXXXXXX}+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\eta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}.$
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\eta\to 0$, and $\kappa\to 0$ in order, we obtain the desired result.
6.2. Large deviations
In this subsection, we study the concentration of the measure $\rho_{\varepsilon}^{Q}$.
Proof of Theorem 1.2.
Note that
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)=\frac{\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]}{Z_{\varepsilon}^{Q}}.$
This implies
$\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)$
$=\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]-\varepsilon\log Z_{\varepsilon}^{Q}.$ (6.21)
Combining (6.21), Proposition 6.1, and Lemma 6.3 yields
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$ (6.22)
This, along with the energy gap estimate in Lemma 3.7, implies
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2}.$ (6.23)
This completes the proof of (1.6).
We now prove (1.8). By following the argument used to obtain (6.22) with Lema 6.3, we get
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-\inf_{\begin{subarray}{c}\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$ (6.24)
Under the condition $\frac{1}{2}e^{-d}\geq c\cdot\delta$, we can apply Lemma 3.10 to obtain
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d}.$ (6.25)
This completes the proof of Theorem 1.2.
Remark 6.4.
By following the proof of Theorem 1.2, we can also establish the following large deviation estimates for the approximating soliton manifolds:
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta\}\big)\leq-c\delta^{2},$ (6.26)
where $\mathcal{M}_{Q}^{\varepsilon}$ is defined as in (4.8), and under $\frac{1}{2}e^{-d}\geq c\cdot\delta$
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d})<\delta\}\big)\leq-ce^{-d},$ (6.27)
where $\mathcal{M}^{\varepsilon,<d}_{Q}$ is defined as in (4.9). Therefore, from (6.26), (6.27) and $\frac{1}{2}e^{-d}\geq c\cdot\delta$, by choosing $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, we have
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta_{\varepsilon}\}\big)\lesssim e^{-c\log\frac{1}{\varepsilon}}$ (6.28)
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}$ (6.29)
for $0<\varepsilon<\varepsilon_{0}$, where $\varepsilon_{0}$ is a small number.
Remark 6.5.
Note that
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d})<\delta\}\subset\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}$ (6.30)
and
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d})<\delta\}$
$\subset\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d})<\delta\big\},$ (6.31)
where
$\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d}$
$:=\Big\{\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j}):-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\;\;\textup{and}\;\;d\leq\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d\Big\}.$
By following the proof of Theorem 1.2 together with Lemma 3.13 and $\frac{1}{2}e^{-d}\geq c\cdot\delta$,
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d})<\delta\big\}\Big)$
$\leq\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d})<\delta\big\}\Big)\leq-c\cdot e^{-(1+\eta)d}.$
Since the distance $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and the collision scale $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, chosen above in (6.29), satisfy $\frac{1}{2}e^{-d_{\varepsilon}}\geq c\cdot\delta_{\varepsilon}$, we have
$\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d})<\delta\big\}\Big)\lesssim\exp\Big\{-\frac{c}{\varepsilon}\cdot e^{-(1+\eta)d_{\varepsilon}}\Big\}\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}.$ (6.32)
Therefore, most of the probability mass is concentrated on the set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}$. For the use of this large deviation estimate, see Remark 7.2 and Lemma 7.1.
7. Preliminaries for the central limit theorem
7.1. Partition into a typical event and large deviation events
Using the large-deviation estimates established above, we decompose the entire space into a typical event and large-deviation events as follows.
Lemma 7.1.
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as in (6.28) and (6.29), and let $F$ be a bounded function. Then
$\int F(\phi)\rho_{\varepsilon}^{Q}(d\phi)=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$
Proof.
Note that, using (6.28), we have
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$
$=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$
$\hphantom{X}+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})\geq\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$ (7.1)
Since $\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}\subset\mathcal{M}_{Q}^{\varepsilon}$
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi).$ (7.2)
Using $\mathcal{M}_{Q}^{\varepsilon}=\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}}\cup\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}}$, we have
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})<\delta_{\varepsilon}\}\cap\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})\geq\delta_{\varepsilon}\}\subset\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}.$
This implies, together with (6.29), that
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}\lesssim\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}.$ (7.3)
Combining (7.1), (7.2), and (7.3) yields
(7.1) $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})+\textup{O}(e^{-c\cdot\varepsilon^{-\frac{1}{2}+}})$
$=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}).$ (7.4)
We further decompose the leading-order term $\mathcal{J}_{1}$ as follows. Using (6.30), we have
$\mathcal{J}_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)$ (7.5)
From the tail bound (6.32), we obtain
$\mathcal{J}_{1}=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\phi)\rho_{\varepsilon}^{Q}(d\phi)+\textup{O}(e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}).$ (7.6)
Combining (7.1), (7.4), and (7.6) completes the proof.
Remark 7.2.
The large deviation estimates (6.28) and (6.29) imply that most of the probability mass is concentrated on the set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq d_{\varepsilon}})<\delta_{\varepsilon}\}$ in (7.2). We note, however, that this set is too large to prove the central limit theorem, since the error term is too large to control on this set. In particular, in (7.13) the first-order term in the Taylor expansion blows up on this set, with size $O(\varepsilon^{0-})$. Therefore, we need to perform a more careful decomposition in (7.5) in order to carry out the Taylor expansion and control the error on the slightly smaller set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}$ in (7.6).
7.2. Coordinate changes
As shown in the previous subsection, we prove that as $\varepsilon\to 0$, most of the probability mass concentrates in the region,
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\},$
where the field $\phi$ is close to the manifold and the solitons are well separated, that is, there is essentially no collision as $d_{\varepsilon}\to\infty$. See the definition of $\mathcal{M}^{\varepsilon,\geq d}_{Q}$ in (4.10).
To prove the central limit theorem (Theorem 1.4), we first perform the following change of coordinates using the disintegration formula in Lemma 4.3.
Remark 7.3.
As discussed in Remark 4.4, the disintegration formula in tangential and normal coordinates has been widely used in settings where the energy functional admits explicit minimizers. However, we emphasize again that, to the best of our knowledge, this is the first time the disintegration formula is applied in a regime where no minimizers exist, and the framework is set up around a multi-soliton manifold.
Proposition 7.4.
Let $Q\in\mathbb{Z}$ with $Q>0$, $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as in (6.28) and (6.29), and let $F$ be a bounded, continuous function. Then
$\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\rho_{\varepsilon}^{Q}(d\phi)$
$=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1}\,\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\xi_{1}\dots,d\xi_{Q},$ (7.7)
where $\pi^{\varepsilon}$ is the projection onto $\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}}$ , defined in (4.15), $\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}$ is the Gaussian measure in Lemma 5.4, and
$U_{\varepsilon}:=\big\{(\xi_{1},\dots,\xi_{Q},v)\in\Delta_{Q}\times V_{\xi_{1},\dots,\xi_{Q}}:\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\;\textup{and}\;\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\big\}$
$\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sin\big(\theta_{x}m_{\xi_{1},\dots,\xi_{Q}}(x)+(1-\theta_{x})\sqrt{\varepsilon}v(x)\big)\cdot(\sqrt{\varepsilon}v)^{3}dx,\quad|\theta_{x}|\leq 1$
$\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v):=\det\big(\operatorname{Id}-W_{\xi_{1},\dots\xi_{Q},\sqrt{\varepsilon}v}\big).$
Here $\Delta_{Q}=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}$. In (7.7), the partition function $\overline{Z}_{\varepsilon}$ represents the integral $\int\cdots\int$ over $U_{\varepsilon}$ without $F$, and $W_{\xi_{1},\dots,\xi_{Q},\sqrt{\varepsilon}v}$ is the Weingarten map, defined in (4.20).
Proof.
Using the disintegration formula in Lemma 4.3 with the coordinate $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ in (4.14), we decompose the integral into its tangential and normal components as follows
$Z[F]:$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}F(\sqrt{\varepsilon}^{-1}(\phi-\pi^{\varepsilon}(\phi)))\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mu_{\varepsilon}^{Q}(d\phi)$
$=\int\cdots\int_{U_{\varepsilon}}F(v)\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v)dx\Big\}$
$\hskip 76.82243pt\cdot\exp\Big\{-\frac{1}{2\varepsilon}\|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}^{2}-\frac{1}{\varepsilon}\langle(-\partial_{x}^{2})m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\sqrt{\varepsilon}v\rangle_{L^{2}}\Big\}$
$\hskip 76.82243pt\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}),$ (7.8)
where $\pi^{\varepsilon}(\phi)=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$, and $\mu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is the Brownian bridge with covariance $(-\partial_{x}^{2})^{-1}$ from 0 to 0, on the normal space, and $d\sigma$ is the surface measure defined in (4.19).
Under the coordinate $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ in (4.14), we perform a Taylor expansion
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos\phi\,dx$ $=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sin(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx$
$\hphantom{X}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})(\sqrt{\varepsilon}v)^{2}dx+\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),$ (7.9)
where $\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}O\big((\sqrt{\varepsilon}v)^{3}\big)dx$. By combining the Cameron–Martin term
$\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx=\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}|^{2}dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(-\partial_{x}^{2}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx+\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}(\sqrt{\varepsilon}v)|^{2}dx$
with the interaction terms (7.9), we obtain
$-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v)$ $=-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle$
$\hphantom{X}-\frac{1}{2\varepsilon}\langle\sqrt{\varepsilon}v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}v\rangle$
$\hphantom{X}+\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),$ (7.10)
where
$E_{L_{\varepsilon}}(\phi)=\frac{1}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}\phi|^{2}dx+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}1-\cos\phi\,dx$
and $\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})=-\partial_{x}^{2}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sin m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}$. On the right-hand side of (7.10), the first two terms and the last term are error terms, whereas the term $\langle v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v)\rangle$ defines the covariance operator of the new Gaussian measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}$. Combining (7.8) and (7.10) yields
$Z[F]$ $=\int\cdots\int_{U_{\varepsilon}}\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle\Big\}$
$\hphantom{XXXXX}\cdot\exp\Big\{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\Big\}\exp\Big\{-\frac{1}{2}\langle v,\cos(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v\rangle\Big\}$
$\hphantom{XXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,\mu^{\perp}_{\xi_{1},\dots,\xi_{Q}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{Q}).$ (7.11)
We first study the first term $\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})$ in (7.11). From Lemma 3.4 and (4.6), we have $E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})=|Q|E_{\textup{kink}}+O(e^{-cd_{\varepsilon}})$, uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$. This implies
$\exp\Big\{-\frac{1}{\varepsilon}E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\Big\}$ $=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}-\frac{1}{\varepsilon}O(e^{-cd_{\varepsilon}})\Big\}$
$=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}(1+O(\varepsilon^{0+}))\Big\},$ (7.12)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}=(1+\eta)\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
We now study the second term $\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle$ in (7.11). By Lemma 3.6 and the condition $\|\sqrt{\varepsilon}v\|_{L^{2}}\leq\delta_{\varepsilon}=\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}$, we have
$\bigg|\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\sqrt{\varepsilon}vdx\bigg|\leq\frac{1}{\varepsilon}\|\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\|_{L^{2}}\|\sqrt{\varepsilon}v\|_{L^{2}}\lesssim\frac{1}{\varepsilon}e^{-(1-\frac{\eta}{2})d_{\varepsilon}}\delta_{\varepsilon}=O(\varepsilon^{0+}),$ (7.13)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}=(1+\eta)\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$. This implies that
$\exp\Big\{-\frac{1}{\varepsilon}\langle\nabla E_{L_{\varepsilon}}(m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),\sqrt{\varepsilon}v\rangle\Big\}=1+O(\varepsilon^{0+})$ (7.14)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$.
We now study the second-order term $\frac{1}{2}\langle v,(-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})v\rangle$ in (7.10), which defines a new base Gaussian measure on the normal space. Note that
$-\partial_{x}^{2}+\cos m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{Q}\text{sech}^{2}(\cdot-\xi_{j})+O(e^{-cd_{\varepsilon}})=\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}+O(e^{-cd_{\varepsilon}}),$
where the error is independent of $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$. Since $v\in V_{\xi_{1},\dots,\xi_{Q}}$ satisfies the orthogonality conditions, from Lemma 5.4, we can define the Gaussian measure
$d\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}=Z_{\xi_{1},\dots,\xi_{Q}}^{-1}\exp\Big\{-\frac{1}{2}\langle v,\mathcal{L}_{\xi_{1},\dots,\xi_{Q}}v\rangle\Big\}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x),$ (7.15)
where $Z_{\xi_{1},\dots,\xi_{Q}}=(Z_{0})^{Q}(1+O(e^{-cd_{\varepsilon}}))$ from Lemma 5.5.
Finally, we study the surface measure $d\sigma(\xi_{1},\dots,\xi_{Q})$ on the multi-soliton manifold. Indeed, since the tangent vectors $\{\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}\}_{j=1}^{Q}$ are almost orthogonal, with an error of order $e^{-cd_{\varepsilon}}$, the Jacobian matrix
$M(\xi_{1},\dots,\xi_{Q})=(\langle\partial_{\xi_{i}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}},\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}\rangle)_{1\leq i,j\leq Q}$
is nearly diagonal, with diagonal entries given by $\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}$, which are independent of $\xi_{1},\dots,\xi_{Q}$. Therefore, we have
$\textup{det}(M(\xi_{1},\dots,\xi_{Q}))=Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}})),$ (7.16)
uniformly in $\xi_{1},\dots\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$.
By using (7.8), (7.11), (7.12), (7.14), (7.15), and (7.16), we can take out terms independent of tangential modes $\xi_{1},\dots,\xi_{Q}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ as follows:
$Z_{\varepsilon}[F]$ $=J_{\varepsilon}\int\cdots\int_{U_{\varepsilon}}F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v)\,d\nu^{\perp}_{m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}}(v)\,d\xi_{Q}\dots d\xi_{1},$
where
$J_{\varepsilon}=\exp\Big\{-\frac{1}{\varepsilon}|Q|E_{\textup{kink}}\Big\}(Z_{0})^{Q}\cdot Q\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd_{\varepsilon}}))(1+O(\varepsilon^{0+})).$
Since the partition function $Z_{\varepsilon}=Z_{\varepsilon}[1]$ contains the same factor $J_{\varepsilon}$, we can cancel the common term $J_{\varepsilon}$ and thus obtain the desired result (7.7).
8. Analysis of the Gaussian measure with Schrödinger operator
In this section, our goal is to establish the correlation decay and the extreme value for the Gaussian measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ on the normal space, with the covariance operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$ (8.1)
subject to Dirichlet boundary conditions, as defined in (5.11).
Remark 8.1.
In this section, we study a joint limit in which $\varepsilon\to 0$ and $L_{\varepsilon}\to\infty$ simultaneously, leading to a competition between energetic and entropic effects. Moreover, in our setting the relevant covariance operator (8.1) is defined around a multi-soliton configuration, which is not a minimizer of the energy. This is in sharp contrast to [19, Theorem 4], where the analysis is based on the covariance operator around an explicit energy minimizer and only the energetic limit $\varepsilon\to 0$ is considered to establish a central limit theorem.
8.1. Correlation decay
In this subsection, we study the decay of correlations for the Gaussian measure with covariance operator $C_{\xi_{1},\dots,\xi_{k}}$ in (8.1). Notice that the operator
$-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
is invertible on the normal space $V_{\xi_{1},\dots,\xi_{k}}$. Therefore, we can take its inverse to study the Green function
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y):=\langle\delta_{x},C_{\xi_{1},\dots,\xi_{k}}\delta_{y}\rangle=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big]$ (8.2)
with Dirichlet boundary conditions on $L_{\varepsilon},L_{\varepsilon}]$. In order to study the decay of correlations in (8.2), we first analyze the projected Ornstein–Uhlenbeck operator on the normal space $V_{\xi_{1},\dots,\xi_{k}}$
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=\big(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big)(x,y)$
with Dirichlet boundary conditions on $L_{\varepsilon},L_{\varepsilon}]$. Note that Dirichlet boundary conditions suppress the variance $(-\partial_{x}^{2}+1)^{-1}(x,x)\approx 0$ near the edges, eliminating randomness there. In particular, the influence of the boundary diminishes exponentially fast as one moves into the interior, and the field in the bulk behaves almost like the infinite-volume Ornstein–Uhlenbeck field. The following proposition shows that the Green function $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
Lemma 8.2.
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
  • (i)
    Bulk regime: Let $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$ and $|\xi_{j}|\leq\overline{L}_{\varepsilon}$, $j=1,\dots,k$, where $\overline{L}_{\varepsilon}$ is defined in (4.2). Then
    $|G_{\textup{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)-ce^{-|x-y|}|\lesssim e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}$
    as $\varepsilon\to 0$, where the exponentially small error term is independent of the tangential modes $\xi_{1},\dots,\xi_{k}$ satisfying $|\xi_{j}|\leq\overline{L}_{\varepsilon}$.
  • (ii)
    Diagonal regime: for any $x\in[-L_{\varepsilon},L_{\varepsilon}]$
    $|G_{\textup{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,x)|\lesssim 1,$
    uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$.
Proof.
We expand the covariance operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$. Recall that the projection operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j}$ (5.12) onto the normal space $V_{\xi_{1},\dots,\xi_{k}}$ is
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ $=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j},$
where each $\mathbf{P}_{j}$ denotes the projection onto the corresponding tangential direction
$\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}.$ (8.3)
Here $t_{1},\dots,t_{k}$ are obtained by applying the Gram–Schmidt orthonormalization to the tangent vectors $\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$, $j=1,\dots,k$. By expanding the projection operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$,
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ $=(-\partial_{x}^{2}+1)^{-1}+\sum_{i,j=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$
$\hphantom{X}-\sum_{j=1}^{k}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$
$\hphantom{X}-\sum_{i=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}$ (8.4)
We study each term separately. We denote by $G^{D}_{\text{OU}}$ the Green function of $(-\partial_{x}^{2}+1)^{-1}$ on $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions. Then, the explicit formula and the bulk behavior are well known as follows:
$G_{\text{OU}}^{D}(x,y)=\frac{\text{sinh}(L_{\varepsilon}-\max\{x,y\})\cdot\text{sinh}(L_{\varepsilon}+\min\{x,y\})}{\text{sinh}(2L_{\varepsilon})}$ (8.5)
and if $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$,
$\big|G_{\text{OU}}^{D}(x,y)-\frac{1}{2}e^{-|x-y|}\big|\lesssim e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}\lesssim e^{-2|\log\varepsilon|}.$ (8.6)
Hence, in the bulk region $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$, the Green function coincides with the whole-line kernel up to an exponentially small error in the distance to the boundary.
We now consider the first projected term $\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$. By direct computation using the definition of $\mathbf{P}_{j}$ in (8.3),
$\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle=t_{i}(x)t_{j}(y)\big\langle t_{i},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle.$ (8.7)
Using Lemma 4.5, that is, $|t_{j}(x)|\lesssim e^{-|x-\xi_{j}|}$, we have
[…]
$\leq e^{-c|x-y|}.$ (8.8)
We next consider the second projection term $(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$ in (8.4). By direct computation using the definition of $\mathbf{P}_{j}$ in (8.3),
$\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}\delta_{y}\big\rangle$ $=t_{j}(y)\big\langle\delta_{x},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle=t_{j}(y)\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz,$ (8.9)
where $G^{D}_{\text{OU}}$ denotes the Green function of $(-\partial_{x}^{2}+1)^{-1}$ on $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions. Note that
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz$ $=\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|x-z|}t_{j}(z)dz+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(G^{D}_{\text{OU}}(x,z)-e^{-|x-z|})t_{j}(z)dz$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}.$ (8.10)
Since $|t_{j}(x)|\lesssim e^{-|x-\xi_{j}|}$, we have
$|\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}|\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|x-z|}e^{-|z-\xi_{j}|}dz\lesssim e^{-c|x-\xi_{j}|}.$ (8.11)
Regarding the error term $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}$, under $|z-\xi_{j}|\leq\frac{1}{2}|\log\varepsilon|$, we use $\max\{|x|,|z|\}\leq\max\{|x|,|\xi_{j}|\}+\frac{1}{2}|\log\varepsilon|$ to obtain
$|\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}|$ $\lesssim\int_{|z-\xi_{j}|\leq\frac{1}{2}|\log\varepsilon|}e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})}t_{j}(z)dz+\int_{|z-\xi_{j}|\geq\frac{1}{2}|\log\varepsilon|}e^{-c|z-\xi_{j}|}dz$
$\lesssim e^{|\log\varepsilon|}e^{-2(L_{\varepsilon}-\max\{|x|,|\xi_{j}|\})}+e^{-c|\log\varepsilon|}$
$\lesssim e^{-|\log\varepsilon|}+e^{-c|\log\varepsilon|},$ (8.12)
where in the last line we used the bulk conditions $|x|\leq L_{\varepsilon}-|\log\varepsilon|$ and $|\xi_{j}|\leq\overline{L}_{\varepsilon}=L_{\varepsilon}-\varepsilon^{-\frac{1}{2}+2\eta}-1$. Hence, by combining (8.10), (8.11), and (8.12), we obtain
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU}}(x,z)t_{j}(z)dz\bigg|\lesssim e^{-c|x-\xi_{j}|}$ (8.13)
as $\varepsilon\to 0$. Then, it follows from (8.9) and (8.13) that
[…]
As for the projected term $\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}$ in (8.4), following the above calculations, we have
$\big|\big\langle\delta_{x},\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\delta_{y}\big\rangle\big|\lesssim e^{-c|x-y|}.$ (8.15)
It follows from (8.4), (8.6), (8.8), (8.14), and (8.15) that
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=ce^{-|x-y|}+O\big(e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}\big)\sim e^{-c|x-y|}$
when $x$ and $y$ are far from the boundaries.
We now prove part (ii) of Lemma 8.2 (diagonal regime). Using the closed form (8.5), we have
$0\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$
for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$. Since $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ is an orthogonal projection, so $0\leq\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\leq\operatorname{Id}$. Hence, adding the projection cannot increase the diagonal
$0\leq G_{\textup{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,x)\leq G^{D}_{\text{OU}}(x,x)\leq\frac{1}{2}$
for all $x\in[-L_{\varepsilon},L_{\varepsilon}]$, uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. This completes the proof of Lemma 8.2.
We are now ready to prove the main part of this subsection. Recall the definition of the Green function corresponding to the covariance operator $\mathcal{C}_{\xi_{1},\dots,\xi_{k}}$ in (8.1)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y):=\langle\delta_{x},C_{\xi_{1},\dots,\xi_{k}}\delta_{y}\rangle=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big].$
The following proposition shows that the Green function $G_{\xi_{1},\dots,\xi_{k}}$ exhibits correlation decay away from the boundary points $L_{\varepsilon}$, $L_{\varepsilon}$.
Proposition 8.3.
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
  • (i)
    Bulk regime: Let $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$ and $|\xi_{j}|\leq\overline{L}_{\varepsilon}$, $j=1,\dots,k$, where $\overline{L}_{\varepsilon}$ is defined in (4.2). Then
    $|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-ce^{-|x-y|}|\lesssim e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})},$
    as $\varepsilon\to 0$, where the exponentially small error term is independent of the tangential modes $\xi_{1},\dots,\xi_{k}$ satisfying $|\xi_{j}|\leq\overline{L}_{\varepsilon}$.
  • (ii)
    Diagonal regime: for any $x\in[-L_{\varepsilon},L_{\varepsilon}]$
    $|G_{\xi_{1},\dots,\xi_{k}}^{D}(x,x)|\lesssim 1,$ (8.16)
    uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$.
Proof.
Using the resolvent identity, we view the Schrödinger operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$
as a perturbation of the Ornstein–Uhlenbeck operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ as follows
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)+(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}WG^{D}_{\xi_{1},\dots,\xi_{k}})(x,y)$
$=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)+(G^{D}_{\xi_{1},\dots,\xi_{k}}WG^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}})(x,y),$ (8.17)
where $W(z)=-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$. From (8.17), we have
$\sum_{j=1}^{k}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$=\sum_{j=1}^{k}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz,$ (8.18)
where $W_{j}(z)=\text{sech}^{2}(z-\xi_{j})$. Note that the equality (8.18) holds after summing over $j$, although the individual components do not commute. Indeed, by Lemma 8.4, we have
$\big\|\big(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}\big)-\big(G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\big)\big\|_{L^{2}\to L^{2}}\lesssim e^{-c\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|},$ (8.19)
where $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$.
Using Lemma 8.2, under the bulk condition $|x|,|y|\leq L_{\varepsilon}-|\log\varepsilon|$ we write
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)$ $=e^{-|x-z|}+O(e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})})$
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)$ $=e^{-|y-z|}+O(e^{-2(L_{\varepsilon}-\max\{|y|,|z|\})}).$ (8.20)
Following the error estimates in (8.10) and (8.12), and using (8.20), $|W_{j}(z)|\lesssim e^{-|z-\xi_{j}|}$, we obtain
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz+e^{-c|\log\varepsilon|}$ (8.21)
for some $\lambda>0$, and
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz+e^{-c|\log\varepsilon|}.$ (8.22)
We proceed by considering two cases: $|y-\xi_{j}|\leq\frac{1}{2}|x-y|$ and $|y-\xi_{j}|\geq\frac{1}{2}|x-y|$. In the former case, we use (8.21), while in the latter we use (8.22).
Case 1: if $|y-\xi_{j}|\leq\frac{1}{2}|x-y|$, then
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\lambda|x-z|}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}(|x-y|-|y-\xi_{j}|-|z-\xi_{j}|)}e^{-\lambda|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz$
$\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz.$ (8.23)
Denoting by $u_{f}=\mathcal{L}^{-1}_{\xi_{1},\dots,\xi_{k}}(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f)$ the solution to the elliptic equation associated with $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$, the elliptic regularity theorem yields
$\|\partial_{x}^{2}u_{f}\|_{L^{2}}^{2}+\|u_{f}\|_{L^{2}}^{2}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}^{2}$ (8.24)
for any $f\in L^{2}$. By using the Sobolev embedding and (8.24), we have that for $f=e^{-\frac{\lambda}{2}|z-\xi_{j}|}$,
$\sup_{y\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-\frac{\lambda}{2}|z-\xi_{j}|}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|$ $\leq C\|u_{f}\|_{H^{\frac{1}{2}+}}\leq C\|u_{f}\|_{H^{2}}\leq C\|\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}f\|_{L^{2}}.$ (8.25)
Combining (8.21), (8.23) and (8.25) yields
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$ (8.26)
Case 2: if $|y-\xi_{j}|\geq\frac{1}{2}|x-y|$, then
$\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\lambda|y-z|}dz$
$\lesssim\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\lambda|z-\xi_{j}|}e^{-\frac{\lambda}{2}(|y-\xi_{j}|-|z-\xi_{j}|)}dz$
$\lesssim e^{-\frac{\lambda}{4}|x-y|}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz.$ (8.27)
Using the elliptic regularity theorem (8.24), and then following (8.25), we obtain
$\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)e^{-\frac{\lambda}{2}|z-\xi_{j}|}dz\bigg|\lesssim 1.$ (8.28)
Combining (8.22), (8.27) and (8.28) yields
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)W_{j}(z)G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(z,y)dz\bigg|\lesssim e^{-\frac{\lambda}{4}|x-y|}.$ (8.29)
We conclude the case study.
We now go back to the resolvent identity (8.17)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)-2\sum_{j=1}^{k}(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}})(x,y).$ (8.30)
If $|y-\xi_{j}|\leq\frac{1}{2}|x-y|$, then we use (8.26) to get
$|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}.$ (8.31)
If $|y-\xi_{j}|\geq\frac{1}{2}|x-y|$, then using the commutator estimate (8.19), we switch the order
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,y)+R_{j}(x,y),$
where
$|R_{j}(x,y)|\lesssim e^{-cd_{\varepsilon}}.$
This, along with (8.29), implies that
$|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)|\lesssim e^{-c|x-y|}+e^{-c|\log(\varepsilon\log\frac{1}{\varepsilon})|}.$ (8.32)
Hence, combining (8.30), (8.31), (8.32), and Lemma 8.2 yields
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=ce^{-|x-y|}+O\big(e^{-2(L_{\varepsilon}-\max\{|x|,|y|\})}\big).$
We now prove the diagonal regime (8.16). From the resolvent identity (8.30) and Lemma 8.2 (ii), we have that for any $x\in[-L_{\varepsilon},L_{\varepsilon}]$
$0\leq G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)\leq G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}(x,x)\lesssim 1,$
uniformly in $\xi_{1},\dots,\xi_{k}$ since $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}$ and $G^{D}_{\xi_{1},\dots,\xi_{k}}$ are positive operators (and hence so are their Green functions) and $W_{j}\geq 0$. This completes the proof of Proposition 8.3.
Before concluding this subsection, we present the proof of the following lemma under the separation condition.
Lemma 8.4.
Under the separation condition $\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|\geq d_{\varepsilon}$, we have
$\big\|\big(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}\big)-\big(G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\big)\big\|_{L^{2}\to L^{2}}\lesssim e^{-c\min_{\ell\neq i}|\xi_{\ell}-\xi_{i}|}\to 0,$
as $\varepsilon\to 0$, where $W_{j}=\textup{sech}^{2}(\cdot-\xi_{j})$.
Proof.
Recall that
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(x,y).$
Using the resolvent identity, we write
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$ $=G^{D}_{\xi_{j}}(x,y)-G^{D}_{\xi_{j}}\Big(\sum_{\ell\neq j}W_{\ell}\Big)G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y),$ (8.33)
where
$G^{D}_{\xi_{j}}(x,y)=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(x,y).$
Again, using the single–bump resolvent identities
$G^{D}_{\xi_{j}}$ $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}+G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{j}}$
$G^{D}_{\xi_{j}}$ $=G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}+G^{D}_{\xi_{j}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}},$
we have
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{j}}=G^{D}_{\xi_{j}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}.$ (8.34)
Inserting the resolvent identity (8.33) into both $G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}$ and $G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}$ and using (8.34), we obtain
$G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}-G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}$
$=\sum_{\ell\neq j}\Big(G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{j}}W_{\ell}G^{D}_{\xi_{1},\dots,\xi_{k}}-G^{D}_{\xi_{1},\dots,\xi_{k}}W_{\ell}G^{D}_{\xi_{j}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\Big)$
$=\sum_{\ell\neq j}(T_{j\ell}+T_{\ell j}).$ (8.35)
By Combes–Thomas bounds for 1D Schrödinger resolvents with a positive mass term, for any bounded $V$ such that the inverse $(-\partial_{x}^{2}+1+V)^{-1}$ exists, we have
$\|\mathbf{1}_{A}(-\partial_{x}^{2}+1+V)^{-1}\mathbf{1}_{B}\|_{L^{2}\to L^{2}}\lesssim e^{-c\cdot\text{dist}(A,B)}.$
This implies that
$\|W_{j}G^{D}_{\xi_{j}}W_{\ell}\|_{L^{2}\to L^{2}}\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}$ (8.36)
since $W_{j}=\text{sech}^{2}(\cdot-\xi_{j})$ is highly localized around $\xi_{j}$ with an exponentially decaying tail. Using (8.36), the operator $T_{j\ell}$ in (8.35) satisfies
$\|T_{j\ell}\|_{L^{2}\to L^{2}}\leq\|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\|W_{j}G^{D}_{\xi_{j}}W_{\ell}\|_{L^{2}\to L^{2}}\|G^{D}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}\lesssim e^{-c|\xi_{j}-\xi_{\ell}|}$ (8.37)
Combining (8.35) and (8.37) yields
$\|G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\xi_{1},\dots,\xi_{k}}-G^{D}_{\xi_{1},\dots,\xi_{k}}W_{j}G^{D}_{\text{OU},\xi_{1},\dots,\xi_{k}}\|_{L^{2}\to L^{2}}$ $\lesssim\sum_{\ell\neq j}\big(\|T_{j\ell}\|_{L^{2}\to L^{2}}+\|T_{\ell j}\|_{L^{2}\to L^{2}}\big)$
$\lesssim\sum_{\ell\neq j}e^{-c|\xi_{j}-\xi_{\ell}|}\lesssim e^{-cd_{\varepsilon}},$
where $d_{\varepsilon}=\min_{\ell\neq j}|\xi_{\ell}-\xi_{j}|\gtrsim|\log(\varepsilon\log\frac{1}{\varepsilon})|$. This completes the proof of Lemma 8.4.
8.2. Maximum of the Gaussian process
In this subsection, we study the maximal behavior of the field $v$ under the Gaussian measure
$d\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}=Z_{\xi_{1},\dots,\xi_{k}}^{-1}e^{-\frac{1}{2}\langle C_{\xi_{1},\dots,\xi_{k}}^{-1}v,v\rangle}\prod_{x\in[-L_{\varepsilon},L_{\varepsilon}]}dv(x),$
defined in Lemma 5.4, with covariance operator $C_{\xi_{1},\dots,\xi_{k}}$ (5.11). In the following proposition, we show that the typical size of $v$ under the measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ satisfies $\|v\|_{L^{\infty}}\leq K\sqrt{|\log L_{\varepsilon}|}$ for some sufficiently large $K\geq 1$.
Lemma 8.5.
There exist a constant $c>0$ such that for sufficiently large $K$, we have
$\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}(A_{\varepsilon}^{c})\lesssim e^{-cK^{2}|\log L_{\varepsilon}|},$ (8.38)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, where
$A_{\varepsilon}:=\{\|v\|_{L^{\infty}([-L_{\varepsilon},L_{\varepsilon}])}\leq K\sqrt{|\log L_{\varepsilon}|}\}.$
Proof.
We first show that there exists a constant $C>0$ such that for any finite $p\geq 2$, we have
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq Cp^{\frac{p}{2}}|x-y|^{\frac{p}{2}},$ (8.39)
uniformly in $(\xi_{1},\dots,\xi_{k})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $\overline{L}_{\varepsilon}$.
Using the fact that $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is a Gaussian measure, we have that for any finite $p\geq 2$
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq p^{\frac{p}{2}}\Big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]\Big)^{\frac{p}{2}}.$ (8.40)
By expanding the second moment, we have
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]$ $=\big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(x)\big]-\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(x)v(y)\big]\big)$
$\hphantom{X}+\big(\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(y)v(y)\big]-\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[v(y)v(x)\big]\big)$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2},$ (8.41)
where
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$ $=G^{D}_{\xi_{1},\dots,\xi_{k}}(x,x)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)$
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}$ $=G^{D}_{\xi_{1},\dots,\xi_{k}}(y,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(y.x).$
Here $G^{D}_{\xi_{1},\dots,\xi_{k}}$ is the Green function defined in (8.2), corresponding to the covariance operator $\mathcal{C}_{\xi_{1},\dots,\xi_{k}}$ given in (5.11). Using the elliptic regularity theorem, one can show that
$\sup_{x\in[-L_{\varepsilon},L_{\varepsilon}]}|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(x,z)|$ $\leq c|y-z|$ (8.42)
$\sup_{y\in[-L_{\varepsilon},L_{\varepsilon}]}|G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)-G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)|$ $\leq c|x-z|,$ (8.43)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and even in $L_{\varepsilon}$. By combining (8.41), (8.42) and (8.43),
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{2}\big]\leq c|x-y|,$ (8.44)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]^{k}$ and $L_{\varepsilon}$. Combining (8.40) and (8.44) yields that for any finite $p\geq 2$,
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|v(x)-v(y)|^{p}\big]\leq cp^{\frac{p}{2}}|x-y|^{\frac{p}{2}},$ (8.45)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $\overline{L}_{\varepsilon}$. Therefore, we obtain the desired estimate (8.39). Once we have (8.39), the maximal behavior of the Gaussian field follows from Talagrand’s chaining argument. See [36, Proposition 5.1] or [37, Chapter 2].
In the previous lemma, we showed that the typical behavior of the field $v$ under the Gaussian measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ is captured by the events $A_{\varepsilon}=\{\|v\|_{L^{\infty}}\leq K\sqrt{|\log L_{\varepsilon}|}\}$. In the following, we study the error estimate outside the typical behavior $A_{\varepsilon}$.
Lemma 8.6.
Let $F$ be a bounded, continuous function and $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$, defined in (6.28). Then
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[F(v)e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\,A_{\varepsilon}^{c},\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]\lesssim e^{-cK^{2}|\log L_{\varepsilon}|},$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, where $\mathcal{E}_{\xi_{1},\dots,\xi_{k}}$ and $\textup{Det}_{\xi_{1},\dots,\xi_{k}}$ are defined in Proposition 7.4.
Proof.
We decompose the event $A_{\varepsilon}^{c}=\{\|v\|_{L^{\infty}}\geq K\sqrt{|\log L_{\varepsilon}|}\}$ into dyadic shells
$\mathbb{E}_{\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}}\Big[e^{\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(\sqrt{\varepsilon}v)^{3}dx}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\,A_{\varepsilon}^{c},\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\sum_{\ell\geq K}\mathbb{E}_{\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}}\Big[e^{\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(\sqrt{\varepsilon}v)^{3}dx}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\,G_{\ell},\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big],$
where $\ell$ ranges over the dyadic numbers $\ell=2^{j}K$, $j\geq 0$, so that the sum runs over $\ell=K,2K,4K,\dots$, and
$G_{\ell}=\Big\{\ell\sqrt{|\log L_{\varepsilon}|}\leq\|v\|_{L^{\infty}}<2\ell\sqrt{|\log L_{\varepsilon}|}\Big\}.$
Since $\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})$ and $\|v\|_{L^{2}}<\eta\sqrt{\log\frac{1}{\varepsilon}}$, using Hölder’s inequality and Lemma 8.5, we obtain
$\sum_{\ell\geq K}\bigg(\mathbb{E}_{\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}}\Big[e^{\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(\sqrt{\varepsilon}v)^{3}dx},\,G_{\ell},\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]\bigg)^{\frac{1}{2}}\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}(G_{\ell})^{\frac{1}{2}}$
$\lesssim\sum_{\ell\geq K}e^{\sqrt{\varepsilon}\log\frac{1}{\varepsilon}\cdot\ell\sqrt{|\log L_{\varepsilon}|}}\cdot(1+\varepsilon^{\frac{k}{2}}\ell^{k}|\log L_{\varepsilon}|^{\frac{k}{2}})\cdot e^{-c\ell^{2}|\log L_{\varepsilon}|}$
$\lesssim\sum_{\ell\geq K}e^{-\frac{c}{2}\ell^{2}|\log L_{\varepsilon}|}\lesssim e^{-cK^{2}|\log L_{\varepsilon}|},$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$ and $L_{\varepsilon}\geq 1$.
9. Ornstein–Uhlenbeck limit on normal space
In this section, we present the central limit theorem on the normal space to the Ornstein–Uhlenbeck measure under the conditions:
$\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$
for every $1\leq j\leq k$. This condition ensures that the test function $g$ is at a distance at least $\varepsilon^{-\frac{1}{2}+2\eta}$ from all solitons $m^{\varepsilon}_{\xi_{1}},\dots,m^{\varepsilon}_{\xi_{k}}$.
Proposition 9.1.
Let $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$ and let $g$ be a real-valued, smooth, compactly supported function. Under the conditions
$\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ (9.1)
for every $1\leq j\leq k$, we have
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$ (9.2)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ and $\overline{L}_{\varepsilon}$. Here $A_{\varepsilon}=\big\{\|v\|_{L^{\infty}}\leq K\sqrt{|\log L_{\varepsilon}|}\big\}$, defined in Lemmas 8.5, and $\delta_{\varepsilon}=\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ in (6.28).
Remark 9.2.
If the tangential modes $\xi_{j}$ do not satisfy the condition $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for every $j=1,\dots,k$ ($g$ is close to one of the solitons), then the central limit theorem to the mean–zero Ornstein–Uhlenbeck measure does not hold. However, in the next section, we show that the contribution of the forbidden region $\textup{dist}(\operatorname*{supp}g,\xi_{j})<\varepsilon^{-\frac{1}{2}+2\eta}$ becomes relatively small compared to the size of the partition function as $L_{\varepsilon}\to 0$. Hence, we can still obtain the desired result with a more careful geometric analysis.
9.1. Asymptotic behavior of the covariance
In order to prove Proposition 9.1, we first take some preliminary steps. With respect to the measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$, we perform the following orthogonal decomposition
$v(x)=\mathbf{P}(x)\langle v,g\rangle+w(x),$ (9.3)
where $\mathbf{P}(x)$ is the projection of $v$ onto the direction of $\langle v,g\rangle$
$\mathbf{P}(x)=\frac{\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[\langle v,g\rangle v(x)\big]}{\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]}.$
The orthogonal decomposition implies that $w$ and $\langle v,g\rangle$ are independent Gaussian random variables. Hence, the measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$ can be decomposed as follows
$d\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}(v)=\frac{1}{\sqrt{2\pi}\sigma_{g}}e^{-\frac{t^{2}}{2\sigma_{g}^{2}}}\mathrm{d}t\,\mathrm{d}\nu^{\perp\perp}_{\xi_{1},\dots,\xi_{k}}(w),$
where the variance $\sigma_{g}^{2}$ is given by
$\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big].$
Before presenting the proof of Proposition 9.1, we study the following lemma, which provides control on the variance $\sigma_{g}^{2}$.
Lemma 9.3.
Let $g$ be a smooth, compactly supported function. Under the conditions
$\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$
for every $1\leq j\leq k$, we have
$\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big),$
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ and $\overline{L}_{\varepsilon}$.
Proof.
By the definition of the Gaussian measure $\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}$, we have
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,G^{D}_{\xi_{1},\dots,\xi_{k}}g\rangle,$ (9.4)
where $G^{D}_{\xi_{1},\dots,\xi_{k}}$ is the Green’s function for the covariance operator
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1+W)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$
defined in (5.11). Here
$W:=-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
is the perturbation term, localized near each soliton center. Using the resolvent identity, we interpret the Schrödinger operator $C_{\xi_{1},\dots,\xi_{k}}$ as a perturbation of the projected Ornstein–Uhlenbeck operator $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ as follows
$G^{D}_{\xi_{1},\dots,\xi_{k}}(x,y)=G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)+(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})(x,y)$ (9.5)
where
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,y)=(\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}})(x,y).$
We first study the leading order term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ in (9.5). By expanding the projection operator $\mathbf{P}_{V_{\xi_{1}\,\dots,\xi_{k}}}$, defined in (5.12), we write
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}$ $=(-\partial_{x}^{2}+1)^{-1}+\sum_{i,j=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$
$\hphantom{X}-\sum_{i=1}^{k}\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}-\sum_{j=1}^{k}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j},$ (9.6)
where $\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}$ is the projection onto the tangential direction. We begin with the first projected term $\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}$. Note that
$\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle=\langle g,t_{i}\rangle\langle g,t_{j}\rangle\big\langle t_{i},(-\partial_{x}^{2}+1)^{-1}t_{j}\big\rangle.$
Under the condition $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$, together with Lemma 4.5 $|t_{j}(x)|\lesssim e^{-|x-\xi_{j}|}$,
$|\langle g,t_{j}\rangle|\lesssim\bigg|\int_{\operatorname*{supp}g}g(x)e^{-|x-\xi_{j}|}dx\bigg|\lesssim\|g\|_{L^{2}}e^{-\textup{dist}(\operatorname*{supp}g,\xi_{j})}\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.7)
This implies that
$|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.8)
Furthermore, the separation condition (9.7) implies
$|\langle g,\mathbf{P}_{i}(-\partial_{x}^{2}+1)^{-1}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}\quad\textup{and}\quad|\langle g,(-\partial_{x}^{2}+1)^{-1}\mathbf{P}_{j}g\rangle|\lesssim e^{-\varepsilon^{-\frac{1}{2}+2\eta}}.$ (9.9)
Combining (9.6), (9.8), and (9.9) yields
$\langle g,G^{D}_{\textup{OU},\xi_{1},\dots,\xi_{k}}g\rangle=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle+O(e^{-\varepsilon^{-\frac{1}{2}+2\eta}}).$ (9.10)
This completes the leading order term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}$ in (9.5).
We now study the perturbation term $G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}}$ in (9.5), which is also an error term. We expand
$|\langle g,(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})g\rangle|$
$\lesssim\sum_{j=1}^{k}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}g(x)\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)e^{-|z-\xi_{j}|}\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)g(y)dy\bigg)dz\bigg)dx\bigg|,$ (9.11)
where we used $\textup{sech}^{2}(z-\xi_{j})\lesssim e^{-|z-\xi_{j}|}$. To use the correlation decay, we separate the Green’s function into its bulk part and boundary parts, as follows, based on Lemma 8.2 and Proposition 8.3,
$G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}(x,z)$ $=ce^{-|x-z|}+O(e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})})$ (9.12)
$G^{D}_{\xi_{1},\dots,\xi_{k}}(z,y)$ $=ce^{-|z-y|}+O(e^{-2(L_{\varepsilon}-\max\{|z|,|y|\})}).$ (9.13)
We first focus on the bulk effect $e^{-|x-z|}$ and $e^{-|z-y|}$ by plugging them into (9.11)
$\sum_{j=1}^{k}\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}g(x)\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|x-z|}e^{-|z-\xi_{j}|}\bigg(\int_{-L_{\varepsilon}}^{L_{\varepsilon}}e^{-|z-y|}g(y)dy\bigg)dz\bigg)dx\bigg|$
$\lesssim\sum_{j=1}^{k}L_{\varepsilon}\bigg|\int_{\operatorname*{supp}g}\int_{\operatorname*{supp}g}e^{-\frac{1}{4}|x-\xi_{j}|}e^{-\frac{1}{4}|y-\xi_{j}|}g(y)g(x)dydx\bigg|$
$\lesssim\sum_{j=1}^{k}L_{\varepsilon}e^{-\frac{1}{2}\text{dist}(\operatorname*{supp}g,\xi_{j})}\|g\|_{L^{1}}\|g\|_{L^{1}}$
$\lesssim L_{\varepsilon}e^{-\frac{1}{2}\varepsilon^{-\frac{1}{2}+2\eta}}\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}},$ (9.14)
where in the last step, we use the condition $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for every $1\leq j\leq k$.
For the boundary effects $e^{-2(L_{\varepsilon}-\max\{|x|,|z|\})}$ and $e^{-2(L_{\varepsilon}-\max\{|z|,|y|\})}$ in (9.12) and (9.13), following the arguments in (8.10) and (8.12), we obtain the error term $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$. Combining (9.11), (9.14), and the boundary error $e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ yields
$|\langle g,(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})g\rangle|\lesssim e^{-\frac{1}{4}\varepsilon^{-\frac{1}{2}+2\eta}}+e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (9.15)
as $\varepsilon\to 0$. Finally, using (9.4), (9.5), (9.10), and (9.15), we obtain
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\big[|\langle v,g\rangle|^{2}\big]$ $=\langle g,G^{D}_{\xi_{1},\dots,\xi_{k}}g\rangle$
$=\langle g,G^{D}_{\textup{OU},\xi_{1},\dots,\xi_{k}}g\rangle+\langle g,(G_{\text{OU},\xi_{1},\dots,\xi_{k}}^{D}WG^{D}_{\xi_{1},\dots,\xi_{k}})g\rangle$
$=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})).$
This completes the proof of Lemma 9.3.
9.2. Ornstein–Uhlenbeck limit on the normal space
We are now ready to prove Proposition 9.1.
Proof of Proposition 9.1.
We decompose the expectation as follows
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\,A_{\varepsilon},\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon},\Big]$
$\hphantom{X}+\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}(e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}-1)\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}+\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2},$ (9.16)
where $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$ is the main term and $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}$ is an error term. On the set $A_{\varepsilon}=\{\|v\|_{L^{\infty}}\leq K\sqrt{|\log L_{\varepsilon}|}\}$, the higher-order error term and the determinant in Proposition 7.4 satisfy
$\Big|\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)\Big|\leq\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}v|^{3}dx$ $\leq\sqrt{\varepsilon}L_{\varepsilon}\|v\|_{L^{\infty}}^{3}$
$\lesssim\varepsilon^{0+}|\log\varepsilon|^{\frac{3}{2}}=\varepsilon^{0+},$
where we used $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, and
$\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)=1+O(\|\sqrt{\varepsilon}v\|_{L^{\infty}}^{k})=1+O(\varepsilon^{\frac{k}{2}-})$ (9.17)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$. This implies that
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{2}=O(\varepsilon^{0+}),$ (9.18)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$.
Regarding the main term $\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}$, we use (9.17) and the tail probability estimates in Lemmas 8.5 and (6.28) to obtain
$\hskip 1.42262pt\text{I}\hskip 1.42262pt_{1}=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}\Big]+O(\varepsilon^{\frac{k}{2}-})+O(e^{-cK^{2}|\log L_{\varepsilon}|})+O(e^{-\frac{c\delta_{\varepsilon}^{2}}{\varepsilon}}),$ (9.19)
where $\delta_{\varepsilon}=\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Using the orthogonal decomposition $v=\mathbf{P}(x)\langle v,g\rangle+w$ in (9.3), where $\langle v,g\rangle$ and $w$ are independent Gaussian, we write
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}\Big]$ $=\iint e^{it}e^{-\frac{t^{2}}{2\sigma_{g}^{2}}}\frac{dt}{\sqrt{2\pi}\sigma_{g}}\nu^{\perp\perp}_{\xi_{1},\dots,\xi_{k}}(w)=e^{-\frac{\sigma_{g}^{2}}{2}}$
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big)\Big\},$ (9.20)
where we used Lemma 9.3
$\sigma_{g}^{2}=\mathbb{E}_{\nu^{\perp}_{m_{\xi_{1},\dots,\xi_{k}}}}\big[|\langle v,g\rangle|^{2}\big]=\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big).$
By combining (9.16), (9.18), (9.19), and (9.20), we obtain
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\cdot\big(1+O(e^{-c\varepsilon^{-\frac{1}{2}+2\eta}})\big)\Big\}+O(\varepsilon^{0+}).$
Therefore, by taking the limit $\varepsilon\to 0$, we have
$\lim_{\varepsilon\to 0}\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{i\langle v,g\rangle}e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]$
$=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}.$
This completes the proof of Proposition 9.1.
Remark 9.4.
Following the proof of Proposition 9.1, we also obtain
$\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{k}}}\Big[e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{k}}(\sqrt{\varepsilon}v),\;A_{\varepsilon},\;\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big]=1+O(\varepsilon^{0+}),$ (9.21)
uniformly in $\xi_{1},\dots,\xi_{k}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$, where $\delta_{\varepsilon}=\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$. Note that in (9.21) we do not need the conditions $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for $j=1,\dots k$. These conditions are only used in (9.20), where we used Lemma 9.3. Note that if $e^{it}$ is replaced by $1$, then (9.20) is immediately equal to $1$.
10. Proof of Theorem 1.4
In this section we prove Theorem 1.4. By studying the geometric structure of the forbidden sets $\textup{dist}(\operatorname*{supp}g,\xi_{j})<\varepsilon^{-\frac{1}{2}+2\eta}$, we remove the conditions used in proving the Ornstein–Uhlenbeck limit (Proposition 9.1): $\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}$ for every $1\leq j\leq k$.
Proof of Theorem 1.4.
Let $g$ be a smooth, compactly supported function. It suffices to prove that
$\lim_{\varepsilon\to 0}\int e^{i\langle\phi,g\rangle}(T^{\varepsilon})_{\#}\rho_{\varepsilon}^{Q}(d\phi)=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\Big\},$
where $T^{\varepsilon}(\phi)=\varepsilon^{-\frac{1}{2}}(\phi-\pi^{\varepsilon}(\phi))$ and $\pi^{\varepsilon}$ denotes the projection onto the multi-soliton manifold $\mathcal{M}^{\varepsilon,\geq(1+\eta)d_{\varepsilon}}_{Q}$, defined in (4.15). Using Lemma 7.1, we decompose the integral into the large-deviation events and the main term $\mathcal{J}_{1}$ as follows
$\int e^{i\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}})$ (10.1)
where
$\mathcal{J}_{1}$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}e^{\langle T^{\varepsilon}(\phi),g\rangle}\rho_{\varepsilon}^{Q}(d\phi),$
Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as defined in (6.28) and (6.29). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.4, we write
$\mathcal{J}_{1}=\overline{Z}_{\varepsilon}^{-1}\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q},$ (10.2)
where
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\psi)=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}}\Big[\psi\cdot e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1}\,\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big].$
Here, the partition function $\overline{Z}_{\varepsilon}$ in (10.2) is
$\overline{Z}_{\varepsilon}=\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}.$ (10.3)
We now split the main term $J_{1}$ in (10.2) into two pats, $\Xi_{\varepsilon}$ and $\Xi_{\varepsilon}^{c}$ as follows
$\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
$\hphantom{X}+\overline{Z}_{\varepsilon}^{-1}\int\cdots\int_{\Xi_{\varepsilon}^{c}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots,d\xi_{Q}$
$=\mathcal{J}_{1}^{(1)}+\mathcal{J}_{1}^{(2)},$ (10.4)
where
$\Xi_{\varepsilon}=\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\textup{dist}(\operatorname*{supp}g,\xi_{j})\geq\varepsilon^{-\frac{1}{2}+2\eta}\;\;\textup{for every $j$}\big\}.$ (10.5)
In order to use Proposition 9.1, we decompose
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})$ $=\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle}\cdot\mathbf{1}_{A_{\varepsilon}})+\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle}\cdot\mathbf{1}_{A_{\varepsilon}^{c}}),$
where $A_{\varepsilon}=\{\|v\|_{L^{\infty}}\leq K\sqrt{|\log L_{\varepsilon}|}\}$. Using the tail estimate Lemma 8.6 with Proposition 9.1, we obtain that on the set $\Xi_{\varepsilon}$
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(e^{i\langle v,g\rangle})=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle_{L^{2}(\mathbb{R})}\Big\}\cdot(1+O(\varepsilon^{0+})),$ (10.6)
uniformly in $(\xi_{1},\dots,\xi_{Q})\in\Xi_{\varepsilon}$. For the term $\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)$ appearing in the partition function $\overline{Z}_{\varepsilon}$ in (10.3), we obtain, using (9.21), that
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)=1+O(\varepsilon^{0+}),$ (10.7)
uniformly in $(\xi_{1},\dots,\xi_{Q})\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]^{k}$.
Combining (10.4), (10.6), (10.3), and (10.7) yields that
$\mathcal{J}_{1}^{(1)}$ $=e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}(\mathbb{R})}}\cdot\frac{\big|\Xi_{\varepsilon}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (10.8)
Furthermore, (10.4), (10.3), and (10.7) imply that
$|\mathcal{J}_{1}^{(2)}|\lesssim\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (10.9)
We show that $\mathcal{J}_{1}^{(1)}\to e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}}}$ and $\mathcal{J}_{1}^{(2)}\to 0$ as $\varepsilon\to 0$. The area of the forbidden region, namely the band of length $\overline{L}_{\varepsilon}^{Q-1}$ and width $d_{\varepsilon}$, is
$\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d_{\varepsilon}\}\big|\sim O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon}).$ (10.10)
Hence, the separation $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ removes only a thin tubular neighbourhood of the diagonal band, of thickness $d_{\varepsilon}$, from the large cube of side length $\overline{L}_{\varepsilon}$, whose volume is of order $O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon})$. Therefore, we have
$\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|$
$=\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\}\big|-O(\overline{L}_{\varepsilon}^{Q-1}d_{\varepsilon})$
$\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q},$ (10.11)
where in the last line we used $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$ and $\overline{L}_{\varepsilon}\sim\varepsilon^{-\frac{1}{2}+\eta}$ from (4.2). So the outside band region has almost full area, up to a relative error of order $\tfrac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\sim\varepsilon^{\frac{1}{2}-}$. For the complement of the set $\Xi_{\varepsilon}$ defined in (10.5), we have
$|\Xi_{\varepsilon}^{c}|$ $\lesssim Q(\overline{L}_{\varepsilon})^{Q-1}\big|\big\{\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]:\textup{dist}(\operatorname*{supp}g,\xi_{j})<\varepsilon^{-\frac{1}{2}+2\eta}\big\}\big|$
$\lesssim Q(\overline{L}_{\varepsilon})^{Q-1}(|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}).$ (10.12)
Combining (10.11) and (10.12) yields that
$\frac{\big|\Xi_{\varepsilon}^{c}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}{\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|}\lesssim\frac{|\operatorname*{supp}g|+2\varepsilon^{-\frac{1}{2}+2\eta}}{\overline{L}_{\varepsilon}}\lesssim\varepsilon^{\eta},$
where we used $\overline{L}_{\varepsilon}=L_{\varepsilon}(1-\varepsilon^{\eta}-\varepsilon^{\frac{1}{2}-\eta})\sim\varepsilon^{-\frac{1}{2}+\eta}$ from (4.2). This implies that in (10.9) $\mathcal{J}_{1}^{(2)}\to 0$ and so in (10.8) $\mathcal{J}_{1}^{(1)}\to e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}}}$ as $\varepsilon\to 0$.
From (10.1), (10.4), and the fact that $\mathcal{J}_{1}^{(1)}\to e^{-\frac{1}{2}\|g\|^{2}_{H^{-1}}}$, $\mathcal{J}_{1}^{(2)}\to 0$ as $\varepsilon\to 0$, we conclude that
$\lim_{\varepsilon\to 0}\int e^{i\langle\phi,g\rangle}(T^{\varepsilon})_{\#}\rho_{\varepsilon}^{Q}(d\phi)=\exp\Big\{-\frac{1}{2}\langle g,(-\partial_{x}^{2}+1)^{-1}g\rangle\Big\}.$
This completes the proof of Theorem 1.4.
11. Proof of Theorem 1.6
Recall that under the coordinate representation $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$, we take the projection $\pi_{\varepsilon}^{T}$ onto the tangential directions by $\pi_{\varepsilon}^{T}(\phi)=(\xi_{1},\dots,\xi_{Q})$. The marginal tangential projection is given by $\pi^{T}_{j}(\phi)=\xi_{j}$, where $\xi_{j}$ denotes the $j$-th ordered center in the increasing rearrangement $\xi_{1}\leq\cdot\leq\xi_{Q}$.
Proof of Theorem 1.6.
Using Lemma 7.1, we decompose the integral into the large-deviation events and the main term $\mathcal{J}_{1}$ as follows
$\int\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi)=\mathcal{J}_{1}+\textup{O}(e^{-c\log\frac{1}{\varepsilon}}),$ (11.1)
where
$\mathcal{J}_{1}$ $=\int_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}}\mathbf{1}_{\{\pi_{\varepsilon}^{T}(\phi)\in A\}}\rho_{\varepsilon}^{Q}(d\phi).$
Here $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, as defined in (6.28) and (6.29). By using the coordinate expression $\phi=m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}+\sqrt{\varepsilon}v$ and Proposition 7.4, we write
$\mathcal{J}_{1}$ $=\overline{Z}_{\varepsilon}^{-1}\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)\mathbf{1}_{\{(\xi_{1},\dots,\xi_{Q})\in A\}}\cdot\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}\,d\xi_{1}\dots d\xi_{Q},$ (11.2)
where
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\psi)=\mathbb{E}_{\nu^{\perp}_{\xi_{1},\dots,\xi_{Q}}}\Big[\psi\cdot e^{\frac{1}{\varepsilon}\mathcal{E}_{\xi_{1}\,\dots,\xi_{Q}}(\sqrt{\varepsilon}v)}\textup{Det}_{\xi_{1},\dots,\xi_{Q}}(\sqrt{\varepsilon}v),\,\|\sqrt{\varepsilon}v\|_{L^{2}}<\delta_{\varepsilon}\Big].$
Using (9.21) and Lemma 8.6, we decompose the main term and the tail contributions as follows
$\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(1)$ $=\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}})+\mathcal{F}_{\xi_{1},\dots,\xi_{Q}}(\mathbf{1}_{A_{\varepsilon}^{c}})$
$=1+O(\varepsilon^{0+}),$ (11.3)
uniformly in $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$. By following the same procedure, we can also write the partition function $\overline{Z}_{\varepsilon}$ in (11.2) as follows
$\overline{Z}_{\varepsilon}=\mathop{\int\cdots\int}\limits_{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}}\mathbf{1}_{\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}}d\xi_{1}\dots d\xi_{Q}\cdot(1+O(\varepsilon^{0+})).$ (11.4)
Combining (11.2), (11.3), and (11.4) yields that
$\mathcal{J}_{1}=\frac{\big|A\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.5)
From (10.11), we have
$\big|\big\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\big\}\cap\{\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}\}\big|$
$\sim(2\overline{L}_{\varepsilon})^{Q}\Big(1-O\Big(\frac{d_{\varepsilon}}{\overline{L}_{\varepsilon}}\Big)\Big)\sim(2\overline{L}_{\varepsilon})^{Q},$ (11.6)
Therefore, the effect of imposing $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq(1+\eta)d_{\varepsilon}$ is negligible compared to the total volume of the simplex $\frac{(2\overline{L}_{\varepsilon})^{Q}}{Q!}$. Using (11.5) and (11.6), we obtain
$\mathcal{J}_{1}=\frac{\big|A\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.7)
Combining (11.1) and (11.7) yields
$\rho_{\varepsilon}^{Q}\{\pi_{\varepsilon}^{T}(\phi)\in A\}=\frac{\big|A\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.8)
Therefore, we obtain the desired result.
By following the arguments used to obtain (11.8), we have
$\rho^{Q}_{\varepsilon}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\frac{\big|\{\xi_{j}\in B\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}<\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}{\big|\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|}(1+O(\varepsilon^{0+})).$ (11.9)
In the following, our goal is to find a density function $f_{j}(x)$
$\rho_{\varepsilon}^{Q}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\int_{B}f_{j}(x)dx\cdot(1+O(\varepsilon^{0+})).$
Fix $x\in(-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon})$ and look at the slice with $\xi_{j}=x$ as follows
$S_{j}(x)=\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\cdots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}:\xi_{j}=x\}.$ (11.10)
Then, $(\xi_{1},\dots,\xi_{j-1})$ form an ordered simplex $\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{j-1}\leq x$ in the interval $\overline{L}_{\varepsilon},x]$, volume
$\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}.$ (11.11)
Also, $(\xi_{j+1},\dots,\xi_{Q})$ form an ordered simplex $x\leq\xi_{j+1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}$ in the interval $x,\overline{L}_{\varepsilon}]$, volume
$\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}.$ (11.12)
Combining (11.10), (11.11) and (11.12) yields
$\big|\{\xi_{j}\in B\}\cap\{-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\}\big|=\int_{B}|S_{j}(x)|dx=\int_{B}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!}dx.$
This, along with (11.9), implies that
$\rho_{\varepsilon}^{Q}\big\{\pi^{T}_{j}(\phi)\in B\big\}=\int_{B}f_{j}(x)dx\cdot(1+O(\varepsilon^{0+})),$
where
$f_{j}(x)=\frac{Q!}{(2L_{\varepsilon})^{Q}}\frac{(x+\overline{L}_{\varepsilon})^{j-1}}{(j-1)!}\frac{(\overline{L}_{\varepsilon}-x)^{Q-j}}{(Q-j)!},\;\;-\overline{L}_{\varepsilon}<x<\overline{L}_{\varepsilon}.$
Recall that $\pi^{T}_{j}(\phi)=\xi_{j}\in[-\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$. Rescaling by $\overline{L}_{\varepsilon}$, define $V_{j}:=\frac{\xi_{j}+\overline{L}_{\varepsilon}}{2\overline{L}_{\varepsilon}}\in[0,1]$. Let $x=2\overline{L}_{\varepsilon}v-\overline{L}_{\varepsilon}$ and $dx=2\overline{L}_{\varepsilon}dv$. Then, the density of $V_{j}$ is
$f_{V_{j}}(v)=f_{j}(x)\cdot(2\overline{L}_{\varepsilon})$ $=\frac{Q!}{(2\overline{L}_{\varepsilon})^{Q}}\frac{(2\overline{L}_{\varepsilon}v)^{j-1}(2\overline{L}_{\varepsilon}(1-v))^{Q-j}}{(j-1)!(Q-j)!}\cdot(2\overline{L}_{\varepsilon})$
$=\frac{Q!}{(j-1)!(Q-j)!}v^{j-1}(1-v)^{Q-j},$
where $0<v<1$. This shows $V_{j}\sim\textup{Beta}(j,Q+1-j)$. Hence, $\pi^{T}_{(j)}(\phi)=\xi_{j}=-\overline{L}_{\varepsilon}+2\overline{L}_{\varepsilon}V_{j}$ follows a Beta distribution, whose expected location is given by
$\mathbb{E}_{\rho_{\varepsilon}^{Q}}\big[\pi^{T}_{j}(\phi)\big]=\Big(-\overline{L}_{\varepsilon}+\frac{2\overline{L}_{\varepsilon}j}{Q+1}\Big)\cdot(1+O(\varepsilon^{0+})).$
This implies that soliton centers $\xi_{1},\dots,\xi_{Q}$ are evenly spaced, dividing the interval $\overline{L}_{\varepsilon},\overline{L}_{\varepsilon}]$ into $Q+1$ equal parts of length $\tfrac{2\overline{L}_{\varepsilon}}{Q+1}$.
Acknowledgements.
H.S. gratefully acknowledges financial support from an NSF grant (CAREER DMS-2044415). The work of P.S. is partially supported by NSF grants DMS-1811093, DMS-2154090, and a Simons Fellowship.